### v2

# Massimiliano Ciranni1 These authors contributed equally to this work.    Vito Paolo Pastore1,2∗    Roberto Di Via1∗    Enzo Tartaglione3    Francesca Odone1    Vittorio Murino2,4

Abstract
Deep learning model effectiveness in classification tasks is often challenged by the quality and quantity of training data which, whenever containing strong spurious correlations between specific attributes and target labels, can result in unrecoverable biases in model predictions. Tackling these biases is crucial in improving model generalization and trust, especially in real-world scenarios. This paper presents Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods in model debiasing while exploiting the inherent bias-learning tendency of diffusion models. Our approach leverages conditional diffusion models to generate synthetic bias-aligned images, used to train a bias amplifier model, to be further employed as an auxiliary method in different unsupervised debiasing approaches. Our proposed method, which also tackles the common issue of training set memorization typical of this type of techniques, beats current state-of-the-art in multiple benchmark datasets by significant margins, demonstrating its potential as a versatile and effective tool for tackling dataset bias in deep learning applications.
footnotetext: Correspondence: vito.paolo.pastore@unige.it
Figure 1: Diffusing DeBias (DDB) Snapshot. A conditional diffusion model learns per-class distributions from a biased dataset, resulting in synthetic generations that reflect these biases (referred to as the bug). These are used to train a Bias Amplifier, converting the bug into a feature useful for debiasing recipes to mitigate bias.
1 Introduction
Deep learning models have shown impressive results in image classification, yet their success depends heavily on training data quality and representativeness. When a deep neural network is trained on a biased dataset, so-called “shortcuts”, corresponding to spurious correlations (i.e., biases) between irrelevant attributes and labels, are learned instead of semantic class attributes, affecting model generalization and impacting test performances. In other words, the model becomes biased towards specific training sub-populations presenting biases, known as bias-aligned, while samples not affected by the bias are referred to as bias-conflicting or unbiased.
In this scenario, several debiasing methods have been proposed [1, 2, 3, 4, 5], addressing the problem under different perspectives. Notably, the majority of available training samples are bias-aligned in a biased dataset (typically, $95\%$ or higher [6]). As such, common supervised debiasing strategies involve reweighting the training set to give more importance to the few available bias-conflicting data [5], for instance, by upsampling [7] or up-weighting [8] them. Conversely, unsupervised debiasing methods, which assume no prior availability of bias information, typically involve an auxiliary model to estimate and convey information on bias for reweighting the training set, employing their signal to guide the training of a debiased target model, while diminishing its tendency to learn bias shortcuts. To this end, there are several existing works [6, 1, 5, 9] exploiting intentionally biased auxiliary models, trained to overfit bias-aligned samples, thus being very confident in both correctly predicting bias-aligned training samples as well as misclassifying bias-conflicting ones [5]. Such a distinct behavior between the two subpopulations is required to convey effective bias supervisory signals for target model debiasing, including gradients or per-sample loss values, which, e.g., should ideally be as high as possible for bias-conflicting and as low as possible for bias-aligned [6]. However, such desired behavior is severely hindered by bias-conflicting training samples, despite their low number. In fact, the low cardinality of this population impedes a suitable modeling of the correct data distribution, yet, they are sufficient to harm the estimation of a reliable bias-aligned data distribution, despite the latter having much higher cardinality. First, bias-conflicting samples in the training set act as noisy labels, affecting the bias learning for the auxiliary models [5]. Furthermore, these models tend to overfit and memorize bias-conflicting training samples within very few epochs, failing to provide different learning signals for the two subpopulations [10]. Despite available strategies, including the usage of the Generalized Cross Entropy (GCE) loss function [6, 3], ensembles of auxiliary models [1, 5], or even ad hoc solutions tailored to the specific dataset considered (e.g., training for only one epoch as in [11]), how to define a protocol for effective and robust auxiliary model training for unsupervised debiasing approaches remains an open issue.
Remarkably, recent works [3, 5, 1] underline how the target model generalization is profoundly impacted by the protocol used to train such an auxiliary model. In principle, if one could remove all bias-conflicting samples from the training set, it would be possible to train a bias-capturing model robust to bias-conflicting samples’ memorization and interference, and capable of conveying ideal learning signals [5].
To this aim, our intuition is to elude these drawbacks by skipping the use of the actual training data in favor of another set of samples synthetically generated by an adequately trained diffusion model. Specifically, we propose to circumvent bias-conflicting interference on auxiliary models leveraging fully synthetic data exploiting intrinsic properties of Conditional Diffusion Probabilistic Models (CDPMs). We aim to learn a per-class biased image distribution, from which we sample an amplified synthetic bias-aligned subpopulation. The resulting synthetic bias-aligned samples are then exploited for training a Bias Amplifier model instead of the original actual training set, solving the issue of bias-conflicting overfitting and interference by construction, as the bias-capturing model does not see the original training set at all. We refer to the resulting approach as Diffusing DeBias (DDB). DDB’s intuition is indirectly supported by recent works studying the problem of biased datasets for image generation tasks  [12, 13], and showing how diffusion models’ generations can be biased, as a consequence of bias present in the training sets (See Sec. 2).
This tendency can be seen as an undesired behavior of generative models, impacting the fairness of generated images, and recent papers have started to propose solutions for decreasing diffusion models’ inclination to learn biases [14, 13, 15]. However, we claim that what appears to be a bug for generative tasks can be turned into a potential feature for image classification model debiasing (see Fig. 1), as it allows the training of a robust auxiliary model that can ideally be plugged into various debiasing schemes. For this reason, we refer to DDB as a plug-in for debiasing methods in image classification. To prove its effectiveness, we incorporate DDB into two different debiasing frameworks, named Recipe I and Recipe II, able to: (i) extract subpopulations further used as pseudo-labels within the popular G-DRO [8] algorithm (Recipe I, Sec. 3.3.1), and (ii) provide a loss function for the training set to be used within an end-to-end method (Recipe II, Sec. 3.3.2). Both approaches outperform the state-of-the-art on popular synthetic and real-world benchmark datasets.
To summarize, our main contributions can be summarized as follows:
  • We introduce DDB, a novel unsupervised debiasing framework that exploits intrinsic properties of CDPMs to learn per-class bias-aligned distributions in biased datasets, with unknown bias information. The resulting CDPM is used for generating synthetic bias-aligned images later employed for training a robust Bias Amplifier model, hence avoiding bias-conflicting training sample memorization and interference (Sec. 3).
  • We propose the usage of DDB as a versatile plug-in for debiasing approaches, designing two recipes: one based on a two-step procedure (Sec. 3.3.1), and the other as an end-to-end (Sec. 3.3.2) algorithm.
  • Both proposed approaches prove to be effective, surpassing state-of-the-art results by a significant margin (Sec. 4) across four different benchmark datasets.
2 Related Work
Various bias mitigation approaches have been explored in the last few years. In the following, we will address the two general setups, supervised and unsupervised, and we finally illustrate how bias is considered in the context of generative diffusion models.
Supervised Debiasing. It refers to methods exploiting available prior knowledge about the bias (e.g., bias attribute annotations), indicating whether a sample exhibits a particular bias (or not). A common strategy consists of training a bias classifier to predict such attributes so that it can guide the target model to extract unbiased representations [16, 17]. Alternatively, bias labels can be thought of as pre-defined encoded subgroups among target classes, allowing for the exploitation of robust training techniques such as G-DRO [8]. At the same time, bias information can be employed to apply regularization schemes aiming at forcing invariance over bias features [18, 19].
Unsupervised Debiasing. It considers bias information unavailable (the most likely scenario in real-world applications), thus having the highest potential impact among the described paradigms. Typical approaches can be distinguished between two-step methods (where the first step identifies bias in training data, generally through pseudo-labeling [11, 9, 4]), and end-to-end (where bias mitigation is performed online with dedicated optimization objectives [6, 20]). An important limitation of current unsupervised methods is avoiding training bias-conflicting overfitting without a properly balanced validation set (a-priori not available), still a major argument of discussion among the community. Employing techniques such as ensembling and noise-robust losses have been proposed, though requiring careful tuning of all the hyperparameters which may require the availability of validation sets with bias annotations to work. In Learning from Failure (LfF), Nam et al. [6] assume that bias features are learned earlier than task-related ones: they employ a bias-capturing model to focus on easier (bias-aligned) samples through the GCE loss [21]. At the same time, a debiased network is jointly trained to assign larger weights to samples that the bias-capturing model struggles to discriminate. Jung et al. [22] introduced a debiasing method where a GAN [23] is trained for style-transfer between several biased appearances. A target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations. In [11], a model is trained with standard ERM and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa: debiasing is brought on by up-sampling the predicted bias-conflicting samples. Notably, they must rely on a validation set with bias annotations to stop the auxiliary model’s training after a few epochs to avoid memorization and to tune the most effective upsampling factor. In [1], a set of bias-amplified classifiers is ensembled (bootstrap ensembling) into a bias-commmittee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample. Finally, a recent trend involves the usage of vision-language models to detect and mitigate biases in classification models [24, 25, 26].
Diffusion Models and Bias. Diffusion models have emerged as powerful generative models capable of producing high-quality data [27]. However, like many deep learning models, they are exposed to biases present in the training data or introduced during the modeling process [15], which can manifest as distorted, unfair, or stereotypical representations deviating from the true underlying data distribution [28]. Several recent studies tackle bias in diffusion models through different approaches, e.g., by exposing and correcting hidden biases in the diffusion process [29], introducing dual-training frameworks for more robust and fair generation [14], addressing specific biases in facial image generation [15], or proposing training methods that counter biased input data effects [13]. While these studies focus on mitigating the bias in diffusion model training aiming to produce fairer and more accurate results, our proposed solution frames this inherent behavior in a novel perspective.
Unlike previously cited works, we show that it is possible to leverage the intrinsic tendency of diffusion models to capture bias in data, using conditioned generations to infer the biased distribution of each target class. This side-effect allows us to sample new synthetic images, sharing the same bias patterns, and use them for unsupervised model debiasing, hence avoiding many limitations of the current state of the art by construction. Actually, exclusively using synthetic data in auxiliary model training prevents overfitting on original real data, while also resulting in a fully-fledged bias amplified model to be exploited for bias identification.
3 Approach
This Section illustrates the proposed debiasing approach, schematically depicted in Figure 2. First, we provide a general description of the problem setting (Sec.  3.1). Second, we report a detailed description of the DDB’s two main components, which include bias diffusion (Sec. 3.2) and the two Recipes for model debiasing (Sec. 3.2).
3.1 Problem Setting
Let us consider a dataset of images and labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$, and assume the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors (e.g., context, appearance, acquisition noise). In this case, it will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors. This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not presenting the same biases. In these conditions, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$. Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or not annotated in a realistic setting [13]. If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization, due to the strong correlation between $b$ and $y$ (often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity [4, 8, 30]), dominating over the true target distribution semantics.
Data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation [12]. For instance, given a CDPM modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we can expect that its generations will be biased [12, 13]. While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature [6]), a CDPM predominantly learns the biased distribution of a specific class, i.e.$\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.
3.2 Diffusing the Bias
In the context of mitigating bias in classification models, the tendency of a CDPM to approximate the per-class biased distribution represents a key feature for training a bias amplification auxiliary model.
Figure 2: Schematic representation of our DDB framework. The debiasing process consists of two key steps: (A) Diffusing the Bias uses a conditional diffusion model with classifier-free guidance to generate synthetic images that preserve training dataset biases, and (B) employs a Bias Amplifier firstly trained on such synthetic data, and subsequently used during inference to extract bias signals from real images. These signals guide debiasing recipes (i.e., two-step and end-to-end methods) that process the original training data to produce a debiased model.
Diffusion Process Background.
The diffusion process progressively converts data into noise through a fixed Markov chain of $T$ steps [31]. Given a data point $\mathbf{x}_{0}$, the forward process adds Gaussian noise according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, resulting in noisy samples $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$. This forward process can be formulated for any timestep $t$ as:  $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ with $\alpha_{s}=1-\beta_{s}$. The reverse process then gradually denoises a sample, reparameterizing each step to predict the noise $\epsilon$ using a model $\boldsymbol{\epsilon}_{\theta}$:
$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$ (1)
where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\sigma_{t}=\sqrt{\beta_{t}}$.
Classifier-Free Guidance for Biased Image Generation.
In cases where additional context or conditioning is available, such as a class label $y$, diffusion models can use this information to guide the reverse process, generating samples that better reflect the target attributes. Classifier-Free Guidance (CFG) [32] introduces a flexible conditioning approach, allowing the model to balance conditional and unconditional outputs without dedicated classifiers.
The CFG technique randomly omits conditioning during training (e.g., with probability $p_{\text{uncond}}=0.1$), enabling the model to learn both generation modalities. During the sampling process, a guidance scale $w$ modulates the influence of conditioning. When $w=0$, the model relies solely on the conditional model. As $w$ increases ($w\geq 1$), the conditioning’s effect is intensified, potentially resulting in more distinct features linked to $y$, thereby increasing fidelity to the class but possibly reducing diversity, whereas lower values help to preserve diversity by decreasing the influence of conditioning. The guided noise prediction is given by:
$\boldsymbol{\epsilon}_{t}=(1+w)\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)-w\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t),$ (2)
where $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)$ is the noise prediction conditioned on class label $y$, and $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction. This modified noise prediction replaces the standard $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ term in the reverse process formula (Equation 1). In this work, we empirically show how the CDPM learns and amplifies underlying biased distribution when trained on a biased dataset with strong spurious correlations, allowing bias-aligned image generation. In the next Section, we describe how such synthetic images can be used to train a robust Bias Amplification model.
3.3 DDB: Bias Amplifier and Model Debiasing
As stated in Sec. 2, a typical unsupervised approach to model debiasing relies on an intentionally biased auxiliary model, named here as Bias Amplifier (BA). This model can be exploited in either two-step or end-to-end approaches, denoted here as Recipe I and Recipe II, respectively.
3.3.1 Recipe I: two-step debiasing
Figure 3: Overview of Recipe I’s two-step debiasing approach.
As a two-step, we adopt the bias supervised algorithm G-DRO [8] as a proven technique for the pure debiasing step. Instead of using the real bias labels (considered unknown), we estimate pseudo-labels performing an inference step by feeding the trained BA with original training data, and identifying as bias-aligned the correctly classified samples, and as bias-conflicting those misclassified. Among possible strategies to assign bias pseudo-labels, such as feature-clustering [9] or anomaly detection [3], we adopt a simple heuristic based on top of the BA misclassifications. Specifically, given a sample $(\mathbf{x}_{i},y_{i},c_{i})$ with $c_{i}$ unknown pseudo-label indicating whether $\mathbf{x}_{i}$ is bias conflicting or aligned, we estimate it as
$\hat{c}_{i}=\mathds{1}\left(\hat{y}_{i}\neq y_{i}~{}\land~{}\mathcal{L}(\hat{y}_{i},y_{i})~{}>~{}\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})\right)$ (3)
where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA model on a real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively. Together with the multiplier $\gamma~{}\in\mathbb{N}$, this condition defines a filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss by a quantity corresponding to a certain z-score of the per-sample training loss distribution (${}{\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})}$ in Equation 3). Once training subgroups pseudo-labels are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 3. This filtering operation refines the plain error set, restricting bias-conflicting selection to the hardest training samples, with potential benefits for the hardest correlation settings. In our experiments, we provide an ablation study comparing filtering and plain error set alternatives, in terms of performance.
3.3.2 Recipe II: end-to-end debiasing
A typical end-to-end debiasing setting includes the joint training of the target debiasing model and one [6] or more [1, 5] intentionally biased auxiliary models. Here, we design an end-to-end debiasing recipe, denoted as Recipe II, incorporating our Bias Amplifier by customizing a widespread general scheme, introduced in Learning from Failure (LfF) [6]. This approach leverages an intentionally biased model trained using the GCE to support the simultaneous training of a debiased model adopting the CE loss re-weighted by a per-sample relative difficulty score. Specifically, we replace the GCE biased model with our Bias Amplifier, which is frozen and only employed in inference to compute its loss function for each original training sample ($\mathcal{L}_{\text{bias\_amp}}$), as schematically represented in Figure 4. Such loss function is used to obtain a weighting factor for the target model’s loss function, defined as $r=\frac{\mathcal{L}_{\text{Bias\_Amp}}}{\mathcal{L}_{\text{debiasing}}+\mathcal{L}_{\text{Bias\_Amp}}}$. Intuitively, $r$ is low for bias-aligned and high for bias-conflicting samples.
Figure 4: Overview of Recipe II’s end-to-end debiasing approach.
4 Experiments
In this Section, after the description of the benchmark datasets employed in our work, we provide general implementation and training details, followed by a discussion of the obtained results and a thorough ablation analysis. In the Suppl. Material, we report Recipe I and Recipe II full implementation details, more examples of generated images, and a quantitative analysis of the dataset bias captured from the CDPM with the assistance of a Vision-Language Model.
Figure 5: Examples of synthetic bias-aligned images across multiple datasets. Each grid shows synthetic images for specific classes across biased datasets, revealing how the model aligns with dataset-specific biases in different contexts.
4.1 Datasets
Our experiments utilize four distinct datasets, a synthetic one, Corrupted CIFAR-10 (CIFAR10-C), and three real-world datasets, namely Waterbirds, Biased Flickr-Faces-HQ (BFFHQ), and Biased Action Recognition (BAR). Specifically, CIFAR10-C [33] is a synthetic dataset derived from CIFAR-10, altered with various corruptions (such as noise, blur, and pixelation): we utilize the configurations with  $\rho=\{0.950,0.980,0.995\}$, the same used in [34]. Waterbirds [35] is an image dataset with inherent biases, exhibiting strong correlations ($\rho=0.950$) between bird species and background environments (e.g., water vs. land). BFFHQ [36] builds upon FFHQ [37] by introducing demographic biases for face recognition  ($y\in\{\text{young, old}\},~{}~{}b\in\{\text{female, male}\},~{}\rho=0.995$). Finally, BAR [6] dataset is crafted to challenge action recognition models that associate specific actions with particular stereotypical contexts.
4.2 Implementation Details
In our experiments, we generate $1000$ synthetic images for each target class using a CDPM implemented as a multi-scale UNet incorporating temporal and conditional embeddings for processing timestep information and class labels. For computational efficiency, input images are resized to $64\times 64$ pixels (except for Corrupted CIFAR-10 whose original size is $32\times 32$, and thus kept as is). Input images are normalized with respect to ImageNet’s mean and standard deviation. The CDPM is trained from scratch for $70,000$, $100,000$, $150,000$, and $200,000$ iterations for the datasets BAR, CIFAR10-C ($\rho=0.950$), Waterbirds, and BFFHQ, respectively. Batch size is set to $32$, and the diffusion process is executed over $1,000$ steps with a linear noise schedule (where $\beta_{1}=1e^{-4}$ and $\beta_{T}=0.028$). Optimization is performed using MSE loss and AdamW optimizer, with an initial learning rate of $1e^{-4}$, adjusted by a CosineAnnealingLR [38] scheduler with a warm-up period for the first $10\%$ of the total iterations. As per the Bias Amplifier training, we use synthetic images from CDPM. A Densenet-121 [39] with ImageNet pre-training is employed across all datasets, featuring a single-layer linear classification head. The training uses regular Cross-Entropy loss function and the AdamW optimizer [40]. For BFFHQ and BAR, the learning rate is $1e^{-4}$, and $5e^{-4}$ for Waterbirds and Corrupted CIFAR-10. Training spans 50 epochs with weight decay $\lambda=0.01$ except for BFFHQ, which uses $\lambda=1.0$ for 100 epochs. We apply standard basic data augmentation strategies (following [4, 6]), including random crops and horizontal flips. For our debiasing recipes we rely on the same backbones used in the existing literature, i.e.ResNet-18 for Corrupted CIFAR-10, BFFHQ, and BAR, and a ResNet-50 for Waterbirds [6, 4, 25]. In Recipe II, $\gamma$ (Equation 3) is set to 3 in all our main experiments. Finally, similarly to [20, 6, 3], we do not rely on a validation set for regularization. This choice is related to real-world application suitability, where no guarantees exist regarding the degree of spurious correlations present in a held-out subset of data used as validation, possibly resulting in detrimental regularization [10].
4.3 Results
4.3.1 Synthetic bias-aligned image generation
First, we present qualitative image generation results across multiple datasets (see Fig. 5), to validate the effectiveness of our bias diffusion approach. The generated samples maintain good fidelity and capture the typical bias patterns of each training set, confirming our conditioned diffusion process’s ability to learn and diffuse inherent biases. This is confirmed by screening the generated images by 2 independent annotators, who were asked to verify (or deny) the presence of the bias in each synthetic sample. In addition, we also carried out an objective evaluation of the generated image content by adopting an image captioner, specifically BLIP-2 [41]. We estimate a caption for each image per class and count the frequency of every word. Suppl. Material (Sec. A.Bias Validation via Captions in Synthetic Images) shows a histogram of the detected keywords related to the Waterbirds dataset, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced. The same behavior also emerges in other dataset used in this study (reported in Sec. A.Bias Validation via Captions in Synthetic Images), providing a good description of the bias attributes. These results highlight our CDPM’s capability to generate synthetic bias-aligned data, setting the stage for our subsequent debiasing techniques.
4.3.2 Debiasing accuracy
Method Corrupted CIFAR-10
$\rho=0.995$ $\rho=0.980$ $\rho=0.950$
[…]
DDB-I (ours) $\bf 31.46\scriptsize{\pm 1.08}$ $\bf 44.90\scriptsize{\pm 1.17}$ $53.38\scriptsize{\pm 0.51}$
[…]
Table 1: Results of DDB Recipes I and II on Corrupted CIFAR-10 and available unsupervised debiasing state-of-the-art works. The best results are highlighted in bold, second-best are underlined.
Tables 1 and 2 report the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets. We report Worst Group Accuracy (WGA) for Waterbirds, as in [8, 5]. For BFFHQ, we report the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following [4, 20]. BAR does not provide bias-attribute annotations, hence we report just the average test accuracy, in line with former works (e.g., [6]). For simplicity, throughout this section, we will refer to Recipe I and Recipe II as DDB-I and DDB-II, respectively. For Corrupted CIFAR-10 (Table 1), DDB-I and DDB-II provide the best performances overall, especially on the most challenging $\rho$ settings, providing a maximum improvement of $1.51\%$ for $\rho=0.995$, $3.12\%$ for $\rho=0.980$, and $4.67\%$ for $\rho=0.950$, all with respect to the DFA method [34]. Notably, both recipes outperform the state-of-the-art performances, showing a stable behavior.
Method Unsup Waterbirds BAR BFFHQ
WGA Avg. Confl.
[…]
LISA [42] $89.20$
[…]
DDB-I
(w/ plain err. set)
[…]
Table 2: Results of DDB Recipes I and II on realistic datasets. Ticks () indicate that the method is bias-unsupervised. We highlight in bold the best results from unsupervised debiasing works, second-best are underlined.
Regarding the considered real-world datasets, we can highlight a similar trend in performance, with DDB’s approaches outperforming the current state of the art (see Table 2). Notably, in Waterbirds, WGA is beating the G-DRO supervised (91.56% vs. 91.40% of G-DRO [8]) and all other unsupervised methods’ accuracies. DDB-I is providing slightly worse performance with respect to DDB-II, but still better than the state of the art. In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN [20]). In BFFHQ, DDB-I and DDB-II obtain accuracies of 74.67% and 70.93%, respectively, as compared with 69.3%, the best state-of-the-art results to date (MoDAD [3]), improving of 6.34 and 2.6 percentage points, respectively.
4.4 Ablation analysis
In this section, we provide an extensive ablation analysis of DDB main components. Without losing generality, all ablations are performed on Waterbirds and exploit Recipe I for model debiasing. Additional ablations on Recipe II can be found in the Suppl. Material. The reported analyses regard several aspects of the data generation process with respect to the Bias Amplifier, the effects of the filtering option in Recipe I, and the behavior of DDB on unbiased data. Biased data generation break point. We aim here to empirically estimate the degree of correlation $\rho$ in the training set leading the CDPM to predominantly learn bias-aligned samples. Starting from the original real images, we build ad hoc controlled training sets such that the number of total samples is kept fixed ($4060$ images, balanced among the classes), while the ratio between aligned and conflicting samples varies. Specifically, we increase the proportion of bias-conflicting from $0.05$ to $0.30$ ($\rho=0.950-0.700$), training a CDPM with CFG on each examined ratio. Then, we generate $1000$ images per class, with two independent annotators manually labeling bias-aligned/conflicting samples. This aims to assess to which extent the bias ratio $\rho$ permits a correct approximation of the biased data distribution, allowing a proper training of the auxiliary model, while investigating the impact on debiasing performance with our Recipe I. The obtained results are highlighted in Table 3. As expected, for $\rho=0.950$, a qualitative analysis reveals very few conflicting generated images. This is still substantially true up to $\rho=0.800$, where we find $8.80\%$ bias-conflicting samples in the generated images, reflected by good debiasing performance ($88.69\%$ and $86.52\%$ for $\rho=0.900$ and $\rho=0.800$). When $\rho=0.700$, as expected, despite the CDPM still amplifying bias-aligned samples, a relevant percentage of bias-conflicting generations emerges ($17.3\%$), impacting final performance.
Input $\rho$ $0.950$ $0.900$ $0.800$ $0.700$
Synthetic output $0.990$ $0.961$ $0.912$ $0.827$
WGA $\bf 90.81$ $88.69$ $86.52$ $67.22$
Table 3: Ablation study on the impact of $\rho$ on correlation bias in synthetic images and WGA performance in Waterbirds.
Bias amplifier data requirements. All our main results are obtained by training a BA on $1000$ synthetic images per class. Here, we want to measure how this choice impacts the effectiveness of our Bias Amplifier. From Table 4 (left), we observe that even 100 synthetic images are enough for good debiasing performance, despite not being competitive with top accuracy scores. As we increase the number of training images, WGA improves, reaching the best performance for $1000$ images per class
CFG strength and generation bias. The most important free parameter for the CDPM with CFG is the guidance conditioning strength $w$. Such strength can affect the variety and consistency of generated images. Here, we investigate DDB dependency on this parameter. Table 4 (right) shows a limited impact on debiasing performance. WGA variations along different $w$ values are limited, with the best performance obtained for $w=1$.
# Synth. images for BA training
# Imgs. $100$ $200$ $500$ $1000$ $2000$
WGA $86.25$ $87.10$ $87.67$ $89.05$ $\bf 90.81$
Guidance strength
$w$ 0 1 2 3 5
WGA 87.85 90.81 89.25 89.56 88.32
Table 4: Ablations on: (left) synthetic image count for training the BA on Waterbirds; (right) Impact of guidance strength $w$ on CDPM sampling, evaluated on Waterbirds, employing Recipe I.
Group extraction via BA plain error sets. In Sec. 3.3.1, we describe how we filter the BA error set through a simple heuristic based on per-sample loss distribution. In Tables 1 and 2, DDB-I (w/ plain err. set, last row in the tables) report our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier. Our results show a trade-off between the two alternatives. The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings ($\rho=0.950$), gaining an advantage from such a coarser selection. On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios ($\rho=0.980$ and $\rho=0.995$).
DDB impact on unbiased dataset. In real-world settings, it is generally unknown whether bias is present in datasets or not. Consequently, an effective approach, while mitigating bias dependency in the case of a biased dataset should also not degrade the performance when the bias is not present. This ablation study aims to assess DDB from this perspective. To obtain this insight, we apply the entire DDB pipeline (adopting Recipe I) on the original CIFAR-10 dataset [44], comparing it with traditional ERM training with CE loss. In this controlled case, Recipe I provides a test accuracy of $84.16\%$, with a slight improvement of $\sim 1\%$ compared to traditional ERM ($83.26\%$) employing the same target model (ResNet18). This result shows that our Bias Amplifier can still provide beneficial signals to the target model in standard learning scenarios, potentially favoring the learning of the most challenging samples.
5 Conclusions
In this work, we present Diffusing DeBias (DDB), a novel debiasing framework capable of learning the per-class bias-aligned training data distribution, leveraging a CDPM. Synthetic images sampled from the inferred biased distribution are used to train a Bias Amplifier, employed to provide a supervisory signal for training a debiased target model. The usage of synthetic bias-aligned data permits avoidance of detrimental effects typically impacting auxiliary models’ training in unsupervised debiasing schemes, such as bias-conflicting sample overfitting and interference, which lead to suboptimal generalization performance of the debiased target model. DDB is effective and versatile, acting as a plug-in for unsupervised debiasing methods. In this work, we design a two-step and an end-to-end debiasing recipes incorporating DDB’s Bias Amplifier: they outperform state-of-the-art works by a margin on four popular biased benchmark datasets, including the most recent relying on vision-language models [25, 45]. Finally, while effectively mitigating bias dependency for biased datasets, DDB does not degrade performance when used with unbiased data, making it suitable for debiasing in real-world applications.
Limitations
DDB’s main limitations rely on the well-known high computational complexity of diffusion models. As a trade-off between computational efficiency and generation quality, we reduce the training sets resolution to $64\times 64$ when training the CDPM. Still, our CDPM requires $\sim 14$ hours for training on the largest considered dataset (i.e., BFFHQ, having $19,200$ images) on an NVIDIA A30 with $24$ GB of VRAM. Another limitation inherited from diffusion models is that the generation quality may decrease if the training dataset is too small, impacting DDB applicability for very small-scale datasets.
Appendix
Ablation Studies with DDB’s Recipe II
This appendix section focuses on additional ablation studies, dedicated to our Recipe II. This set of ablations aims at providing additional evidence of Recipe II’s robustness and effectiveness, similar to what is shown in Sec. 4.4 of the main paper regarding Recipe I. As in the main paper, we run all our ablations on Waterbirds.
Biased data generation break point.
Using the same protocol employed for Recipe I, we want to study the impact of the input training set for the CDPM on the whole pipeline. In the main paper, we describe the process of building ad-hoc training sets (where all have a total of $4060$ images, balanced among the classes, but a different ratio between aligned and conflicting samples) from Waterbirds, and the subsequent generation of $1000$ images per class, with two independent annotators manually labeling bias-aligned/conflicting samples. Using the same pools of generated images, we train a dedicated Bias Amplifier and use it in our Recipe II. The obtained results, outlined in Table 5, resemble what is shown in the main paper for Recipe I (Table 3). As expected, when the input training set for the CDPM has $\rho=0.900$, resulting generations are still substantially dominated by aligned samples, allowing for training an effective Bias Amplifier. This is reflected by the final WGA ($90.03\%$). A similar behavior can be found for $\rho=0.800$ (WGA = $89.53\%$). As it was for Recipe I, $\rho=0.700$ results in a sensibly less effective Bias Amplifier due to the relevant percentage of conflicting samples among generated images. As a consequence, the final WGA is negatively affected ($73.52\%$).
Input $\rho$ $0.950$ $0.900$ $0.800$ $0.700$
Synthetic output $0.990$ $0.961$ $0.912$ $0.827$
WGA $\bf 91.56$ $90.03$ $89.53$ $73.52\%$
Table 5: Ablation on the number of training images for the Bias Amplifier, in terms of resulting WGA on Waterbirds, when using DDB’s Recipe II.
Bias amplifier data requirements.
# Synth. images for BA training
# Imgs. $100$ $200$ $500$ $1000$ $2000$
WGA $84.42$ $90.03$ $90.18$ $90.78$ $\bf 91.56$
Table 6: Ablation on the number of training images for the Bias Amplifier, in terms of WGA on Waterbirds, when using DDB’s Recipe II.
This ablation study measures how the number of training images for the Bias Amplifier impacts DDB’s Recipe II. Results are evaluated in terms of WGA on Waterbirds (see Table 4). $200$ images ($100$ images per class) are enough to obtain state-of-the-art competitive results (WGA equal to $90.03$), with an increasing number of images corresponding to higher WGA values, up to a maximum of $91.56$ for $2000$ total images.
CFG strength and generation bias.
Guidance Strength
$w$ 00 $\bf 1$ $2$ $3$ $5$
WGA $87.07$ $\bf 91.56$ $88.01$ $89.87$ $88.30$
Table 7: Impact of guidance strength $w$ on the final WGA on Waterbirds, when using DDB’s Recipe II.
Here, we evaluate DDB’s Recipe II dependency on the guidance strength parameter. Results are summarized in Table 7. The best results correspond to the usage of $w=1$, with different guidance strength parameter values having a relatively limited impact on debiasing performance.
DDB impact on an unbiased dataset.
Similarly to what is shown in Sec. 4.4, we want to ensure that applying Recipe II on an unbiased dataset does not degrade the final performance. Also Recipe II demonstrates to be applicable in unbiased scenarios, as it measures a final test accuracy of $83.87\%$ on the original test set of CIFAR10, which is slightly higher than the ERM baseline ($83.26\%$).
Full Implementation Details of Recipe I (two-step)
For the hyperparameters of Recipe I, we use for each dataset the following configurations. Corrupted CIFAR-10: regardless of the specific $\rho$, we set batch size = 32, learning rate = $1\times e^{-4}$, weight decay = 0.01, and training is set to a total of 150 epochs. BAR: batch size = 32, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80. Waterbirds: batch size = 128, learning rate = $5\times e^{-4}$, weight decay = 1.0, training for 60 epochs. BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 1.0, training for 60 epochs. The GDRO optimization is always run with the --robust and --reweight_groups flags.
Full Implementation Details of Recipe II (end-to-end)
For Corrupted CIFAR-10 and BAR, we use the same parameters reported in [6]. Corrupted CIFAR-10: Regardless of the specific $\rho$, we set batch size = 256, learning rate = $5\times e^{-4}$, weight decay = $0.01$, and training is set to a total of 50 epochs. BAR: batch size = 128, learning rate = $1\times e^{-4}$, weight decay = 0.01, training epochs = 50. Waterbirds: batch size = 128, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80. BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 0.01, training for 50 epochs. For the real-world datasets, we accumulate gradients for $8$ iterations ($16$ for BFFHQ) with mini-batches made of $16$ samples.
Additional Biased Generations
In this Section, we present $100$ additional per class generated images for each dataset utilized in the study, comprising three real-world and one synthetic dataset, as produced by our CDPM. With CIFAR10-C (Figure 6), the model proficiently synthesizes classes with specific corruptions, evidencing its ability to replicate synthetic bias patterns. In the Waterbirds dataset (Figure 7), the model adeptly captures and replicates the correlation between bird species and their background. In the BFFHQ dataset (Figure 8), it effectively reflects demographic biases by mirroring gender-appearance correlations. The results from the BAR dataset (Figure 9) exhibit the model’s adeptness in handling multiple bias patterns, preserving typical environmental contexts for different actions. Overall, the model generates high-quality samples that are fidelity-biased across all datasets. Our research confirms that the model exhibits bias-learning behavior, as it not only learns but also amplifies existing biases within diverse datasets. We leverage this characteristic to produce synthetic biased data aimed at enhancing the robustness of debiasing techniques.
Figure 6: Synthetic bias-aligned image generations for each class of the CIFAR10-C dataset. Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
Figure 7: Synthetic bias-aligned image generations for each class of the Waterbirds dataset. Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
Figure 8: Synthetic bias-aligned image generations for each class of the BFFHQ dataset. Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
Figure 9: Synthetic bias-aligned image generations for each class of the BAR dataset. Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
Bias Validation via Captions in Synthetic Images
As an ulterior confirmation of bias in CDPM-generated images, besides two independent annotators, we implement an unsupervised analysis pipeline using the pre-trained BLIP-2 FlanT5XL model [41] for zero-shot image captioning without any task-specific prompting, for avoiding any guidance in the descriptions of the synthetic data. Our method processes the resulting captions through stop word removal and frequency analysis to reveal underlying biases without relying on predetermined categories or supervised classification models. Figure 10 presents the keyword frequency distribution for the Waterbirds dataset, where the most prevalent terms naturally correspond to known bias attributes (e.g., environmental contexts). The same is true for the keyword frequency distribution of the BFFHQ dataset (Figure 11) and BAR (Figure 12) where the most prevalent terms are related to the gender and the context respectively. These findings confirm that generated images are bias-aligned. Furthermore, the obtained word frequency histograms suggest that images generated by DDB’s bias diffusion step may provide information useful to support bias identification (or discovery) at the dataset level.
Figure 10: Word frequency analysis from 1000 generated captions for Waterbirds classes ‘landbird’ (left) and ‘waterbird’ (right), showing top 10 most frequent terms.
Figure 11: Word frequency analysis from 1000 generated captions for BFFHQ classes ‘young’ (left) and ‘old’ (right), showing top 10 most frequent terms.
Figure 12: Word frequency analysis from 1000 generated captions for BAR classes, showing top 10 most frequent terms for each class.
(a) Class 0
(b) Class 1
(c) Class 2
(d) Class 3
(e) Class 4
(f) Class 5

### v3

@@
-Massimiliano Ciranni1 These authors contributed equally to this work.    Vito Paolo Pastore1,2∗    Roberto Di Via1∗    Enzo Tartaglione3    Francesca Odone1    Vittorio Murino2,4
+Diffusing DeBias: Synthetic Bias Amplification for Model Debiasing
 Abstract
-Deep learning model effectiveness in classification tasks is often challenged by the quality and quantity of training data which, whenever containing strong spurious correlations between specific attributes and target labels, can result in unrecoverable biases in model predictions.
-Tackling these biases is crucial in improving model generalization and trust, especially in real-world scenarios.
-This paper presents Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods in model debiasing while exploiting the inherent bias-learning tendency of diffusion models.
-Our approach leverages conditional diffusion models to generate synthetic bias-aligned images, used to train a bias amplifier model, to be further employed as an auxiliary method in different unsupervised debiasing approaches.
-Our proposed method, which also tackles the common issue of training set memorization typical of this type of techniques, beats current state-of-the-art in multiple benchmark datasets by significant margins, demonstrating its potential as a versatile and effective tool for tackling dataset bias in deep learning applications.
+Deep learning model effectiveness in classification tasks is often challenged by the quality and quantity of training data whenever they are affected by strong spurious correlations between specific attributes and target labels.
+This results in a form of bias affecting training data, which typically leads to unrecoverable weak generalization in prediction.
+This paper aims at facing this problem by leveraging bias amplification with generated synthetic data: we introduce Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods of unsupervised model debiasing exploiting the inherent bias-learning tendency of diffusion models in data generation.
+Specifically, our approach adopts conditional diffusion models to generate synthetic bias-aligned images, which replace the original training set for learning an effective bias amplifier model that we subsequently incorporate into an end-to-end and a two-step unsupervised debiasing approach.
+By tackling the fundamental issue of bias-conflicting training samples memorization in learning auxiliary models, typical of this type of techniques, our proposed method beats current state-of-the-art in multiple benchmark datasets, demonstrating its potential as a versatile and effective tool for tackling bias in deep learning models.
 footnotetext: Correspondence: vito.paolo.pastore@unige.it
-Figure 1: Diffusing DeBias (DDB) Snapshot.
-A conditional diffusion model learns per-class distributions from a biased dataset, resulting in synthetic generations that reflect these biases (referred to as the bug).
-These are used to train a Bias Amplifier, converting the bug into a feature useful for debiasing recipes to mitigate bias.
+Figure 1: Comparison of bias amplifier training approaches.
+(a) Traditional methods that use real biased datasets lead to the memorization of biased-conflicting samples, resulting in suboptimal auxiliary models.
+(b) Our proposed approach leverages conditional diffusion models to learn class-specific biases and amplify them into synthetic images.
+Such generations can replace the original training set for learning an effective Bias Amplifier, eliminating memorization effects by avoiding exposure to real data.
 1 Introduction
@@
 To this aim, our intuition is to elude these drawbacks by skipping the use of the actual training data in favor of another set of samples synthetically generated by an adequately trained diffusion model.
-Specifically, we propose to circumvent bias-conflicting interference on auxiliary models leveraging fully synthetic data exploiting intrinsic properties of Conditional Diffusion Probabilistic Models (CDPMs).
+Specifically, we propose to circumvent bias-conflicting interference on auxiliary models leveraging fully biased synthetic data exploiting intrinsic properties of Conditional Diffusion Probabilistic Models (CDPMs).
 We aim to learn a per-class biased image distribution, from which we sample an amplified synthetic bias-aligned subpopulation.
@@
 This tendency can be seen as an undesired behavior of generative models, impacting the fairness of generated images, and recent papers have started to propose solutions for decreasing diffusion models’ inclination to learn biases [14, 13, 15].
-However, we claim that what appears to be a bug for generative tasks can be turned into a potential feature for image classification model debiasing (see Fig.
-1), as it allows the training of a robust auxiliary model that can ideally be plugged into various debiasing schemes.
+However, we claim that what is considered a problem for generative tasks can be turned into a feature for image classification model debiasing, as it allows the training of a robust auxiliary model that can ideally be plugged into various debiasing schemes, solving bias-conflicting training samples memorization, by construction (see Figure 1).
 For this reason, we refer to DDB as a plug-in for debiasing methods in image classification.
-To prove its effectiveness, we incorporate DDB into two different debiasing frameworks, named Recipe I and Recipe II, able to: (i) extract subpopulations further used as pseudo-labels within the popular G-DRO [8] algorithm (Recipe I, Sec.
+To prove its effectiveness and versatility, we incorporate DDB’s bias amplifier into two different debiasing frameworks, named Recipe I and Recipe II, able to: (i) extract subpopulations further used as pseudo-labels within the popular G-DRO [8] algorithm (Recipe I, Sec.
 3.3.1), and (ii) provide a loss function for the training set to be used within an end-to-end method (Recipe II, Sec.
 3.3.2).
-Both approaches outperform the state-of-the-art on popular synthetic and real-world benchmark datasets.
+Both approaches outperform the state-of-the-art on popular benchmark datasets.
 To summarize, our main contributions can be summarized as follows:
-  • We introduce DDB, a novel unsupervised debiasing framework that exploits intrinsic properties of CDPMs to learn per-class bias-aligned distributions in biased datasets, with unknown bias information.
+  • We introduce DDB, a novel and effective unsupervised debiasing framework that exploits intrinsic properties of CDPMs to learn per-class bias-aligned distributions in biased datasets, with unknown bias information.
 The resulting CDPM is used for generating synthetic bias-aligned images later employed for training a robust Bias Amplifier model, hence avoiding bias-conflicting training sample memorization and interference (Sec.
@@
 2 Related Work
-Various bias mitigation approaches have been explored in the last few years.
-In the following, we will address the two general setups, supervised and unsupervised, and we finally illustrate how bias is considered in the context of generative diffusion models.
+In this section, we provide a broad categorization of model debasing works.
 Supervised Debiasing.
-It refers to methods exploiting available prior knowledge about the bias (e.g., bias attribute annotations), indicating whether a sample exhibits a particular bias (or not).
+Supervised methods exploit available prior knowledge about the bias (e.g., bias attribute annotations), indicating whether a sample exhibits a particular bias (or not).
 A common strategy consists of training a bias classifier to predict such attributes so that it can guide the target model to extract unbiased representations [16, 17].
@@
 It considers bias information unavailable (the most likely scenario in real-world applications), thus having the highest potential impact among the described paradigms.
-Typical approaches can be distinguished between two-step methods (where the first step identifies bias in training data, generally through pseudo-labeling [11, 9, 4]), and end-to-end (where bias mitigation is performed online with dedicated optimization objectives [6, 20]).
-An important limitation of current unsupervised methods is avoiding training bias-conflicting overfitting without a properly balanced validation set (a-priori not available), still a major argument of discussion among the community.
-Employing techniques such as ensembling and noise-robust losses have been proposed, though requiring careful tuning of all the hyperparameters which may require the availability of validation sets with bias annotations to work.
+Among such methods, a popular strategy consists in involving auxiliary models at any level of the debiasing process.
+We categorize these methods into two-step  [11, 9, 4]), and end-to-end  [6, 20]), in the following paragraphs.
+End-to-end approaches.
+Here, bias mitigation is performed online with dedicated optimization objectives.
+Employing techniques such as ensembling and noise-robust losses have been proposed, though they require careful tuning of all the hyperparameters, which may require the availability of validation sets with bias annotations to work.
 In Learning from Failure (LfF), Nam et al.
@@
 At the same time, a debiased network is jointly trained to assign larger weights to samples that the bias-capturing model struggles to discriminate.
-Jung et al.
-[22] introduced a debiasing method where a GAN [23] is trained for style-transfer between several biased appearances.
-A target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations.
-In [11], a model is trained with standard ERM and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa: debiasing is brought on by up-sampling the predicted bias-conflicting samples.
+In DebiAN, Li et al.
+[20] jointly train an auxiliary model (the discoverer) and a classifier (the target model) in an alternate scheme so that the discoverer can identify bias learned by the classifier during training, optimizing an ad-hoc loss function for bias mitigation.
+Two-step methods.
+Here, the auxiliary model is employed in a first step to identify bias in training data, generally through pseudo-labeling, while the second stage consists in actual bias mitigation exploiting the inferred pseudo-labels.
+In [11], the auxiliary model is trained with standard ERM and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa: debiasing is brought on by up-sampling the predicted bias-conflicting samples.
 Notably, they must rely on a validation set with bias annotations to stop the auxiliary model’s training after a few epochs to avoid memorization and to tune the most effective upsampling factor.
-In [1], a set of bias-amplified classifiers is ensembled (bootstrap ensembling) into a bias-commmittee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample.
-Finally, a recent trend involves the usage of vision-language models to detect and mitigate biases in classification models [24, 25, 26].
-Diffusion Models and Bias.
-Diffusion models have emerged as powerful generative models capable of producing high-quality data [27].
-However, like many deep learning models, they are exposed to biases present in the training data or introduced during the modeling process [15], which can manifest as distorted, unfair, or stereotypical representations deviating from the true underlying data distribution [28].
-Several recent studies tackle bias in diffusion models through different approaches, e.g., by exposing and correcting hidden biases in the diffusion process [29], introducing dual-training frameworks for more robust and fair generation [14], addressing specific biases in facial image generation [15], or proposing training methods that counter biased input data effects [13].
-While these studies focus on mitigating the bias in diffusion model training aiming to produce fairer and more accurate results, our proposed solution frames this inherent behavior in a novel perspective.
-Unlike previously cited works, we show that it is possible to leverage the intrinsic tendency of diffusion models to capture bias in data, using conditioned generations to infer the biased distribution of each target class.
-This side-effect allows us to sample new synthetic images, sharing the same bias patterns, and use them for unsupervised model debiasing, hence avoiding many limitations of the current state of the art by construction.
-Actually, exclusively using synthetic data in auxiliary model training prevents overfitting on original real data, while also resulting in a fully-fledged bias amplified model to be exploited for bias identification.
-3 Approach
-This Section illustrates the proposed debiasing approach, schematically depicted in Figure 2.
-First, we provide a general description of the problem setting (Sec.
-3.1).
-Second, we report a detailed description of the DDB’s two main components, which include bias diffusion (Sec.
+In [1], a set of auxiliary classifiers is ensembled (bootstrap ensembling) into a bias-commmittee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample.
+Finally, a recent trend involves the usage of vision-language models to identify bias, further exploiting bias predictions to mitigate biases in classification models [22, 23, 24].
+Among the described methods, a subset exploit the auxiliary model for bias amplification ([5, 6, 9]) under the general assumption that a bias amplification model would either misclassify or be less confident on bias-conflicting samples, with final debiasing performance strongly dependent on such behavior, and specifically, on how strongly the auxiliary model captures training bias [3, 5].
+However, this fundamental assumption only holds if bias-conflicting samples are not memorized during training, which is likely to happen in very few training epochs [10, 3], especially considering that the very few training bias-conflicting samples are more likely to be overfitted, as reported in [8, 25].
+One could argue that an annotated validation set could be used for properly regularizing the bias amplification model, but the latter cannot be assumed available a priori in real cases, this being en emerging argument of discussion among the community [10, 3].
+For this reason, in this work and differently from the previously cited research, we propose that syntethic bias-aligned samples could be used instead of the original training set, solving the bias-conflicting memorization issue by construction.
+However, an unsupervised real-world scenario poses a problem for generating a pure distribution of bias-aligned samples.
+As a solution, we propose to leverage diffusion models for generating a synthetic bias-aligned training set to be used for training a bias amplifier, inspired by recent works, showing how they can manifest distorted, unfair, or stereotypical representations deviating from the true underlying data distribution [26], when exposed to biases present in the training data  [15].
+While the vast majority of available works regarding generative models and debiasing refer to strategies for obtaining unbiased synthetic data generation [27, 28], up to our knowledge, we are the first to propose the usage of generative models for bias amplification.
+The closest available work, employing generative models for model debiasing, is Jung et al.
+[29], where the authors introduce a debiasing method exploiting a GAN [30] trained for style-transfer between several biased appearances, with a different objective.
+Here, a target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations.
+In our work, we show that it is possible to leverage the intrinsic tendency of diffusion models to capture bias in data, using conditioned generations to infer the biased distribution of each target class, thus allowing the sampling of new synthetic images, sharing the same bias pattern of original train data.
+Our method provides a memorization-free bias amplification model, capable of providing precise supervisory signals for accurate model debiasing virtually compatible with any debiasing approach involving auxiliary models.
+Figure 2: Schematic representation of our DDB framework.
+The debiasing process consists of two key steps: (A) Diffusing the Bias uses a conditional diffusion model with classifier-free guidance to generate synthetic images that preserve training dataset biases, and (B) employs a Bias Amplifier firstly trained on such synthetic data, and subsequently used during inference to extract supervisory bias signals from real images.
+These signals are used to guide the training process of a target debiased model by designing two debiasing recipes (i.e., 2-step and end-to-end methods).
+3 The Approach
+Our proposed debiasing approach is schematically depicted in Figure 2.
+In this section, we provide at first a general description of the problem setting (Sec.
+3.1), and then, we illustrate in detail DDB’s two main components, which include bias diffusion (Sec.
 3.2) and the two Recipes for model debiasing (Sec.
-3.2).
+3.3).
 3.1 Problem Setting
-Let us consider a dataset of images and labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$, and assume the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors (e.g., context, appearance, acquisition noise).
-In this case, it will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors.
+Let us consider a general data distribution $p_{\text{data}}$, typically encompassing multiple factors of variation and classes, and to build a dataset of images with the associated labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$ sampled from such a distribution.
+Let us also assume that the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors of variations, i.e.possible biases such as context, appearance, acquisition noise, viewpoint, etc..
+In this case, data will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors.
 This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not presenting the same biases.
-In these conditions, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$.
-Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or not annotated in a realistic setting [13].
-If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization, due to the strong correlation between $b$ and $y$ (often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity [4, 8, 30]), dominating over the true target distribution semantics.
-Data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation [12].
-For instance, given a CDPM modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we can expect that its generations will be biased [12, 13].
-While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature [6]), a CDPM predominantly learns the biased distribution of a specific class, i.e.$\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.
+In the same way, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$.
+Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or merely not annotated, in a realistic setting [13] In this context, biased and unbiased samples equivalently refer to bias-aligned and bias conflicting samples..
+If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization.
+This is due to the strong correlation between $b$ and $y$, often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity [4, 8, 31]), which is dominating over the true target distribution semantics.
+It is important to notice that data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation [12].
+For instance, given a Conditional Diffusion Probabilistic Models (CDPM) modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ (with parameters $\phi$) that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we expect that its generations will be biased, as also stated in [12, 13].
+While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature [6]), a CDPM predominantly learns the biased distribution of a specific class, i.e., $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.
 3.2 Diffusing the Bias
-In the context of mitigating bias in classification models, the tendency of a CDPM to approximate the per-class biased distribution represents a key feature for training a bias amplification auxiliary model.
-Figure 2: Schematic representation of our DDB framework.
-The debiasing process consists of two key steps: (A) Diffusing the Bias uses a conditional diffusion model with classifier-free guidance to generate synthetic images that preserve training dataset biases, and (B) employs a Bias Amplifier firstly trained on such synthetic data, and subsequently used during inference to extract bias signals from real images.
-These signals guide debiasing recipes (i.e., two-step and end-to-end methods) that process the original training data to produce a debiased model.
-Diffusion Process Background.
+In the context of mitigating bias in classification models, the tendency of a CDPM to approximate the per-class biased distribution represents a key feature for training an auxiliary bias amplified model.
+The Diffusion Process.
 The diffusion process progressively converts data into noise through a fixed Markov chain of $T$ steps [32].
@@
 This forward process can be formulated for any timestep $t$ as:  $q(\mathbf{x}_{t}|\mathbf{x}_{0})=\mathcal{N}(\mathbf{x}_{t};\sqrt{\bar{\alpha}_{t}}\mathbf{x}_{0},(1-\bar{\alpha}_{t})\mathbf{I})$, where $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ with $\alpha_{s}=1-\beta_{s}$.
-The reverse process then gradually denoises a sample, reparameterizing each step to predict the noise $\epsilon$ using a model $\boldsymbol{\epsilon}_{\theta}$:
-$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$ (1)
+The reverse process then gradually denoises a sample, reparameterizing each step to predict the noise $\epsilon$ using a model $\bm{\epsilon}_{\theta}$:
+$\mathbf{x}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\left(\mathbf{x}_{t}-\frac{\beta_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)\right)+\sigma_{t}\mathbf{z},$ (1)
 where $\mathbf{z}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$ and $\sigma_{t}=\sqrt{\beta_{t}}$.
 Classifier-Free Guidance for Biased Image Generation.
-In cases where additional context or conditioning is available, such as a class label $y$, diffusion models can use this information to guide the reverse process, generating samples that better reflect the target attributes.
+In cases where additional context or conditioning is available, such as a class label $y$, diffusion models can use this information to guide the reverse process, generating samples that better reflect the target attributes and semantics.
 Classifier-Free Guidance (CFG) [33] introduces a flexible conditioning approach, allowing the model to balance conditional and unconditional outputs without dedicated classifiers.
@@
 When $w=0$, the model relies solely on the conditional model.
-As $w$ increases ($w\geq 1$), the conditioning’s effect is intensified, potentially resulting in more distinct features linked to $y$, thereby increasing fidelity to the class but possibly reducing diversity, whereas lower values help to preserve diversity by decreasing the influence of conditioning.
+As $w$ increases ($w\geq 1$), the conditioning effect is intensified, potentially resulting in more distinct features linked to $y$, thereby increasing fidelity to the class while possibly reducing diversity, whereas lower values help to preserve diversity by decreasing the influence of conditioning.
 The guided noise prediction is given by:
-$\boldsymbol{\epsilon}_{t}=(1+w)\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)-w\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t),$ (2)
-where $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)$ is the noise prediction conditioned on class label $y$, and $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction.
-This modified noise prediction replaces the standard $\boldsymbol{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ term in the reverse process formula (Equation 1).
-In this work, we empirically show how the CDPM learns and amplifies underlying biased distribution when trained on a biased dataset with strong spurious correlations, allowing bias-aligned image generation.
-In the next Section, we describe how such synthetic images can be used to train a robust Bias Amplification model.
+$\bm{\epsilon}_{t}=(1+w)\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)-w\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t),$ (2)
+where $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t,y)$ is the noise prediction conditioned on class label $y$, and $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ is the unconditional noise prediction.
+This modified noise prediction replaces the standard $\bm{\epsilon}_{\theta}(\mathbf{x}_{t},t)$ term in the reverse process formula (Equation 1).
+In this work, we empirically show how CDPM learns and amplifies the underlying biased distribution when trained on a biased dataset with strong spurious correlations, allowing bias-aligned image generation.
 3.3 DDB: Bias Amplifier and Model Debiasing
 As stated in Sec.
-2, a typical unsupervised approach to model debiasing relies on an intentionally biased auxiliary model, named here as Bias Amplifier (BA).
-This model can be exploited in either two-step or end-to-end approaches, denoted here as Recipe I and Recipe II, respectively.
-3.3.1 Recipe I: two-step debiasing
-Figure 3: Overview of Recipe I’s two-step debiasing approach.
-As a two-step, we adopt the bias supervised algorithm G-DRO [8] as a proven technique for the pure debiasing step.
-Instead of using the real bias labels (considered unknown), we estimate pseudo-labels performing an inference step by feeding the trained BA with original training data, and identifying as bias-aligned the correctly classified samples, and as bias-conflicting those misclassified.
-Among possible strategies to assign bias pseudo-labels, such as feature-clustering [9] or anomaly detection [3], we adopt a simple heuristic based on top of the BA misclassifications.
-Specifically, given a sample $(\mathbf{x}_{i},y_{i},c_{i})$ with $c_{i}$ unknown pseudo-label indicating whether $\mathbf{x}_{i}$ is bias conflicting or aligned, we estimate it as
+2, a typical unsupervised approach to model debiasing relies on an auxiliary intentionally-biased model, named here as Bias Amplifier (BA).
+This model can be exploited in either 2-step or end-to-end approaches, denoted here as Recipe I and Recipe II, respectively.
+3.3.1 Recipe I: 2-step debiasing
+Figure 3: Overview of Recipe I’s 2-step debiasing approach.
+The adopted 2-step approach consists in 1) applying the auxiliary model trained on biased generated data to perform a bias pseudo-labeling, hence estimating bias-aligned/bias-conflict split of original actual data, and 2) apply a bias supervised method to train a debiased target model for classification.
+For the latter, we use the group DRO algorithm [8] (G-DRO) as a proven technique for the pure debiasing step.
+In other words, being in the unsupervised bias scenario where the real bias labels are unknown, we estimate bias pseudo-labels performing an inference step by feeding the trained BA with the original actual training data, and identifying as bias-aligned the correctly classified samples, and as bias-conflicting those misclassified.
+Among possible strategies to assign bias pseudo-labels, such as feature-clustering [9] or anomaly detection [3], we adopt a simple heuristic based on the BA misclassifications.
+Specifically, given a sample $(\mathbf{x}_{i},y_{i},c_{i})$ with $c_{i}$ unknown pseudo-label indicating whether $\mathbf{x}_{i}$ is bias conflicting or aligned, we estimate bias-conflicting samples as
 $\hat{c}_{i}=\mathds{1}\left(\hat{y}_{i}\neq y_{i}~{}\land~{}\mathcal{L}(\hat{y}_{i},y_{i})~{}>~{}\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})\right)$ (3)
-where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA model on a real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively.
-Together with the multiplier $\gamma~{}\in\mathbb{N}$, this condition defines a filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss by a quantity corresponding to a certain z-score of the per-sample training loss distribution (${}{\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})}$ in Equation 3).
-Once training subgroups pseudo-labels are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 3.
-This filtering operation refines the plain error set, restricting bias-conflicting selection to the hardest training samples, with potential benefits for the hardest correlation settings.
-In our experiments, we provide an ablation study comparing filtering and plain error set alternatives, in terms of performance.
+where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA on the real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively, depending on the loss $\mathcal{L}$.
+Together with the multiplier $\gamma~{}\in\mathbb{N}$, this condition defines a sort of filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss increased by a quantity corresponding to a certain z-score of the per-sample training loss distribution (${}{\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})}$ in Eq.
+3).
+Once bias pseudo-labels over original training data are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 3.
+The above filtering operation refines the plain error set, restricting bias-conflicting sample selection to the hardest training samples, with potential benefits for the most difficult correlation settings ($\rho>0.99$).
+Later in the experimental section, we provide an ablation study comparing different filtering ($\gamma$) configurations and plain error set alternatives.
 3.3.2 Recipe II: end-to-end debiasing
-A typical end-to-end debiasing setting includes the joint training of the target debiasing model and one [6] or more [1, 5] intentionally biased auxiliary models.
-Here, we design an end-to-end debiasing recipe, denoted as Recipe II, incorporating our Bias Amplifier by customizing a widespread general scheme, introduced in Learning from Failure (LfF) [6].
-This approach leverages an intentionally biased model trained using the GCE to support the simultaneous training of a debiased model adopting the CE loss re-weighted by a per-sample relative difficulty score.
+A typical end-to-end debiasing setting includes the joint training of the target debiasing model and one [6] or more [1, 5] auxiliary intentionally-biased models.
+Here, we design an end-to-end debiasing procedure, denoted as Recipe II, incorporating our BA by customizing a widespread general scheme, introduced in the Learning from Failure (LfF) method [6].
+LfF leverages an intentionally-biased model trained using Generalized CE (GCE) loss to support the simultaneous training of a debiased model adopting the CE loss re-weighted by a per-sample relative difficulty score.
 Specifically, we replace the GCE biased model with our Bias Amplifier, which is frozen and only employed in inference to compute its loss function for each original training sample ($\mathcal{L}_{\text{bias\_amp}}$), as schematically represented in Figure 4.
-Such loss function is used to obtain a weighting factor for the target model’s loss function, defined as $r=\frac{\mathcal{L}_{\text{Bias\_Amp}}}{\mathcal{L}_{\text{debiasing}}+\mathcal{L}_{\text{Bias\_Amp}}}$.
-Intuitively, $r$ is low for bias-aligned and high for bias-conflicting samples.
+Such loss function is used to obtain a weighting factor for the target model loss function, defined as $r=\frac{\mathcal{L}_{\text{Bias\_Amp}}}{\mathcal{L}_{\text{debiasing}}+\mathcal{L}_{\text{Bias\_Amp}}}$.
+Coarsely speaking, $r$ should be low for bias-aligned and high for bias-conflicting samples.
 Figure 4: Overview of Recipe II’s end-to-end debiasing approach.
@@
 In this Section, after the description of the benchmark datasets employed in our work, we provide general implementation and training details, followed by a discussion of the obtained results and a thorough ablation analysis.
-In the Suppl.
-Material, we report Recipe I and Recipe II full implementation details, more examples of generated images, and a quantitative analysis of the dataset bias captured from the CDPM with the assistance of a Vision-Language Model.
+In the Supplementary Material, we report Recipe I and Recipe II full implementation details, more examples of generated images, and a quantitative analysis of the dataset bias captured from the CDPM with the assistance of a Vision-Language Model.
 Figure 5: Examples of synthetic bias-aligned images across multiple datasets.
@@
 4.1 Datasets
-Our experiments utilize four distinct datasets, a synthetic one, Corrupted CIFAR-10 (CIFAR10-C), and three real-world datasets, namely Waterbirds, Biased Flickr-Faces-HQ (BFFHQ), and Biased Action Recognition (BAR).
-Specifically, CIFAR10-C [33] is a synthetic dataset derived from CIFAR-10, altered with various corruptions (such as noise, blur, and pixelation): we utilize the configurations with  $\rho=\{0.950,0.980,0.995\}$, the same used in [34].
-Waterbirds [35] is an image dataset with inherent biases, exhibiting strong correlations ($\rho=0.950$) between bird species and background environments (e.g., water vs.
+Our experiments utilize five distinct datasets: Waterbirds, Biased Flickr-Faces-HQ (BFFHQ), Biased Action Recognition (BAR), and ImageNet-9/ImageNet-A.
+Waterbirds  [8] is an image dataset exhibiting strong correlations ($\rho=0.950$) between bird species and background environments (e.g., water vs.
 land).
-BFFHQ [36] builds upon FFHQ [37] by introducing demographic biases for face recognition  ($y\in\{\text{young, old}\},~{}~{}b\in\{\text{female, male}\},~{}\rho=0.995$).
-Finally, BAR [6] dataset is crafted to challenge action recognition models that associate specific actions with particular stereotypical contexts.
+It has 4,795 training images, 1,199 images for validation, and 5,794 for testing.
+BFFHQ [4] builds upon FFHQ [34] by introducing demographic biases for face recognition, while BAR  [6] is crafted to challenge action recognition models that associate specific actions with particular stereotypical contexts.
+For these two last datasets, we specifically refer to their original versions introduced respectively in [6] and [4]: BFFHQ has a total of 19,200 training images where  ( $y\in\{\text{young, old}\},~{}~{}b\in\{\text{female, male}\},~{}\rho=0.995$).
+Then, it provides 1,000 validation images (all bias-aligned) and 1,000 testing images with uniformly distributed labels and bias attributes.
+BAR presents 6 different classes, but does not provide explicit bias annotations or a validation set, with 1,941 images for training and 654 for testing.
+Finally, ImageNet-A consists of 7,500 real-world images from a 200-class subset of ImageNet, where deep learning models consistently struggle to make correct predictions.
+Originally introduced by Hendrycks et al.
+[35] to evaluate adversarial robustness, it is also employed as a model debiasing benchmark, usually as a bias-conflicting evaluation set for models trained on ImageNet-9 [36], a collection of images with 57,200 samples where categories are super-classes of ImageNet, put together to exhibit textural and shape biases [36].
+As such, both our CDPM and Bias Amplifier are trained on ImageNet-9.
+Then, we evaluate both recipes on ImageNet-A in terms of Average Accuracy, following the same protocol of [36, 1, 31]
 4.2 Implementation Details
 In our experiments, we generate $1000$ synthetic images for each target class using a CDPM implemented as a multi-scale UNet incorporating temporal and conditional embeddings for processing timestep information and class labels.
-For computational efficiency, input images are resized to $64\times 64$ pixels (except for Corrupted CIFAR-10 whose original size is $32\times 32$, and thus kept as is).
+For computational efficiency, input images are resized to $64\times 64$ pixels.
 Input images are normalized with respect to ImageNet’s mean and standard deviation.
-The CDPM is trained from scratch for $70,000$, $100,000$, $150,000$, and $200,000$ iterations for the datasets BAR, CIFAR10-C ($\rho=0.950$), Waterbirds, and BFFHQ, respectively.
+The CDPM is trained from scratch for $70,000$, $150,000$, $200,000$, $100.000$ iterations for the datasets BAR, Waterbirds, BFFHQ, and ImageNet-9 respectively.
 Batch size is set to $32$, and the diffusion process is executed over $1,000$ steps with a linear noise schedule (where $\beta_{1}=1e^{-4}$ and $\beta_{T}=0.028$).
@@
 The training uses regular Cross-Entropy loss function and the AdamW optimizer [39].
-For BFFHQ and BAR, the learning rate is $1e^{-4}$, and $5e^{-4}$ for Waterbirds and Corrupted CIFAR-10.
+For BFFHQ and BAR, the learning rate is $1e^{-4}$, and $5e^{-4}$ for Waterbirds, and ImageNet-9.
 Training spans 50 epochs with weight decay $\lambda=0.01$ except for BFFHQ, which uses $\lambda=1.0$ for 100 epochs.
 We apply standard basic data augmentation strategies (following [4, 6]), including random crops and horizontal flips.
-For our debiasing recipes we rely on the same backbones used in the existing literature, i.e.ResNet-18 for Corrupted CIFAR-10, BFFHQ, and BAR, and a ResNet-50 for Waterbirds [6, 4, 25].
+For our debiasing recipes, we rely on the same backbones used in the existing literature, i.e.ResNet-18 for BFFHQ, BAR, ImageNet-9/A, and a ResNet-50 for Waterbirds [6, 4, 23].
 In Recipe II, $\gamma$ (Equation 3) is set to 3 in all our main experiments.
@@
 5), to validate the effectiveness of our bias diffusion approach.
-The generated samples maintain good fidelity and capture the typical bias patterns of each training set, confirming our conditioned diffusion process’s ability to learn and diffuse inherent biases.
+The generated samples maintain good fidelity and capture the typical bias patterns of each training set per class, confirming our conditioned diffusion process’s ability to learn and diffuse inherent biases.
 This is confirmed by screening the generated images by 2 independent annotators, who were asked to verify (or deny) the presence of the bias in each synthetic sample.
@@
 Material (Sec.
-A.Bias Validation via Captions in Synthetic Images) shows a histogram of the detected keywords related to the Waterbirds dataset, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced.
-The same behavior also emerges in other dataset used in this study (reported in Sec.
-A.Bias Validation via Captions in Synthetic Images), providing a good description of the bias attributes.
+A.E) shows a histogram of the detected keywords related to the Waterbirds dataset, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced.
+The same behavior also emerges in other datasets used in this study, where a consistent description of the bias attributes can be noted, as reported in Sec.
+A.E.
 These results highlight our CDPM’s capability to generate synthetic bias-aligned data, setting the stage for our subsequent debiasing techniques.
-4.3.2 Debiasing accuracy
-Method Corrupted CIFAR-10
-$\rho=0.995$ $\rho=0.980$ $\rho=0.950$
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-DDB-I (ours) $\bf 31.46\scriptsize{\pm 1.08}$ $\bf 44.90\scriptsize{\pm 1.17}$ $53.38\scriptsize{\pm 0.51}$
-[…]
-Table 1: Results of DDB Recipes I and II on Corrupted CIFAR-10 and available unsupervised debiasing state-of-the-art works. The best results are highlighted in bold, second-best are underlined.
-Tables 1 and 2 report the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets.
+4.3.2 Classification accuracy
+Table 1 reports the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets.
 We report Worst Group Accuracy (WGA) for Waterbirds, as in [8, 5].
-For BFFHQ, we report the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following [4, 20].
-BAR does not provide bias-attribute annotations, hence we report just the average test accuracy, in line with former works (e.g., [6]).
+For BFFHQ, we report the average and the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following [4, 20].
+BAR and ImageNet-A do not provide bias-attribute annotations, hence, we report just the average test accuracy.
 For simplicity, throughout this section, we will refer to Recipe I and Recipe II as DDB-I and DDB-II, respectively.
-For Corrupted CIFAR-10 (Table 1), DDB-I and DDB-II provide the best performances overall, especially on the most challenging $\rho$ settings, providing a maximum improvement of $1.51\%$ for $\rho=0.995$, $3.12\%$ for $\rho=0.980$, and $4.67\%$ for $\rho=0.950$, all with respect to the DFA method [34].
-Notably, both recipes outperform the state-of-the-art performances, showing a stable behavior.
-Method Unsup Waterbirds BAR BFFHQ
-WGA Avg. Confl.
+It’s worth noticing that, for BAR and BFFHQ, we report benchmark results related to works that employ the same versions as ours.
+Method Unsup Waterbirds BAR BFFHQ ImageNet-A
+WGA Avg. Avg. Confl. Avg.
 […]
@@
 […]
+Park et al. [44] $71.68$
+[…]
+[…]
+CDvG+LfF [29] $84.80$ $62.20\text{\scriptsize{$\pm$0.45}}$ $34.60$
+[…]
 […]
@@
 […]
-Table 2: Results of DDB Recipes I and II on realistic datasets.
+Table 1: Results of DDB Recipes I and II on benchmark datasets.
 Ticks () indicate that the method is bias-unsupervised.
 We highlight in bold the best results from unsupervised debiasing works, second-best are underlined.
-Regarding the considered real-world datasets, we can highlight a similar trend in performance, with DDB’s approaches outperforming the current state of the art (see Table 2).
-Notably, in Waterbirds, WGA is beating the G-DRO supervised (91.56% vs.
-91.40% of G-DRO [8]) and all other unsupervised methods’ accuracies.
+Across all the considered benchmark datasets, both DDB recipes show state-of-the-art results (see Table 1).
+Notably, in Waterbirds, DDB’s WGA is higher than G-DRO supervised (91.56% vs.
+91.40% of G-DRO [8]) and all the other unsupervised methods’ ones.
 DDB-I is providing slightly worse performance with respect to DDB-II, but still better than the state of the art.
-In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN [20]).
-In BFFHQ, DDB-I and DDB-II obtain accuracies of 74.67% and 70.93%, respectively, as compared with 69.3%, the best state-of-the-art results to date (MoDAD [3]), improving of 6.34 and 2.6 percentage points, respectively.
+In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN [20]), in terms of conflicting accuracy.
+In BFFHQ, our DDB-I reaches an accuracy of 74.67%, with an average improvement of 1% when compared to ETF-Debias [43] (73.60%), the best state-of-the-art results to date, and provides a boost of 7.45% over the recent DeNetDM [45].
 4.4 Ablation analysis
@@
 The reported analyses regard several aspects of the data generation process with respect to the Bias Amplifier, the effects of the filtering option in Recipe I, and the behavior of DDB on unbiased data.
-Biased data generation break point.
-We aim here to empirically estimate the degree of correlation $\rho$ in the training set leading the CDPM to predominantly learn bias-aligned samples.
-Starting from the original real images, we build ad hoc controlled training sets such that the number of total samples is kept fixed ($4060$ images, balanced among the classes), while the ratio between aligned and conflicting samples varies.
-Specifically, we increase the proportion of bias-conflicting from $0.05$ to $0.30$ ($\rho=0.950-0.700$), training a CDPM with CFG on each examined ratio.
-Then, we generate $1000$ images per class, with two independent annotators manually labeling bias-aligned/conflicting samples.
-This aims to assess to which extent the bias ratio $\rho$ permits a correct approximation of the biased data distribution, allowing a proper training of the auxiliary model, while investigating the impact on debiasing performance with our Recipe I.
-The obtained results are highlighted in Table 3.
-As expected, for $\rho=0.950$, a qualitative analysis reveals very few conflicting generated images.
-This is still substantially true up to $\rho=0.800$, where we find $8.80\%$ bias-conflicting samples in the generated images, reflected by good debiasing performance ($88.69\%$ and $86.52\%$ for $\rho=0.900$ and $\rho=0.800$).
-When $\rho=0.700$, as expected, despite the CDPM still amplifying bias-aligned samples, a relevant percentage of bias-conflicting generations emerges ($17.3\%$), impacting final performance.
-Input $\rho$ $0.950$ $0.900$ $0.800$ $0.700$
-Synthetic output $0.990$ $0.961$ $0.912$ $0.827$
-WGA $\bf 90.81$ $88.69$ $86.52$ $67.22$
-Table 3: Ablation study on the impact of $\rho$ on correlation bias in synthetic images and WGA performance in Waterbirds.
 Bias amplifier data requirements.
@@
 Here, we want to measure how this choice impacts the effectiveness of our Bias Amplifier.
-From Table 4 (left), we observe that even 100 synthetic images are enough for good debiasing performance, despite not being competitive with top accuracy scores.
-As we increase the number of training images, WGA improves, reaching the best performance for $1000$ images per class
+From Table 2 (left), we observe that even 100 synthetic images are enough for good debiasing performance despite not being competitive with top accuracy scores.
+As we increase the number of training images, WGA improves, reaching the best performance for $2000$ images per class.
 CFG strength and generation bias.
@@
 3.3.1, we describe how we filter the BA error set through a simple heuristic based on per-sample loss distribution.
-In Tables 1 and 2, DDB-I (w/ plain err.
-set, last row in the tables) report our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier.
+In Table 1, the last row (DDB-I w/ plain err.
+set) reports our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier.
 Our results show a trade-off between the two alternatives.
-The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings ($\rho=0.950$), gaining an advantage from such a coarser selection.
-On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios ($\rho=0.980$ and $\rho=0.995$).
+The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings ($\rho=0.950$) (e.g., Waterbirds), gaining an advantage from such a coarser selection.
+On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios, like BFFHQ ($\rho=0.995$).
+Bias-Identification Accuracy To frame the goodness of our Bias Amplifier in correctly identifying aligned and conflicting training samples, we provide the accuracy of such process when considering samples misclassified by the BA.
+Specifically, we report a bias-identification accuracy of $89.00\%$ for Waterbirds and $96.00\%$ on BFFHQ.
+Our very high accuracy supports the effectiveness of using the proposed protocol, which is not dependent from any careful regularization, bias-annotated validation sets, or large ensembles of auxiliary classifiers.
 DDB impact on unbiased dataset.
@@
 This ablation study aims to assess DDB from this perspective.
-To obtain this insight, we apply the entire DDB pipeline (adopting Recipe I) on the original CIFAR-10 dataset [44], comparing it with traditional ERM training with CE loss.
+To obtain this insight, we apply the entire DDB pipeline (adopting Recipe I) on a common image dataset such as CIFAR-10  [46], comparing it with traditional ERM training with CE loss.
 In this controlled case, Recipe I provides a test accuracy of $84.16\%$, with a slight improvement of $\sim 1\%$ compared to traditional ERM ($83.26\%$) employing the same target model (ResNet18).
@@
 Finally, while effectively mitigating bias dependency for biased datasets, DDB does not degrade performance when used with unbiased data, making it suitable for debiasing in real-world applications.
-Limitations
+Limitations.
 DDB’s main limitations rely on the well-known high computational complexity of diffusion models.
@@
 Appendix
-Ablation Studies with DDB’s Recipe II
+Appendix A Ablation Studies with DDB’s Recipe II
 This appendix section focuses on additional ablation studies, dedicated to our Recipe II.
@@
 As in the main paper, we run all our ablations on Waterbirds.
-Biased data generation break point.
-Using the same protocol employed for Recipe I, we want to study the impact of the input training set for the CDPM on the whole pipeline.
-In the main paper, we describe the process of building ad-hoc training sets (where all have a total of $4060$ images, balanced among the classes, but a different ratio between aligned and conflicting samples) from Waterbirds, and the subsequent generation of $1000$ images per class, with two independent annotators manually labeling bias-aligned/conflicting samples.
-Using the same pools of generated images, we train a dedicated Bias Amplifier and use it in our Recipe II.
-The obtained results, outlined in Table 5, resemble what is shown in the main paper for Recipe I (Table 3).
-As expected, when the input training set for the CDPM has $\rho=0.900$, resulting generations are still substantially dominated by aligned samples, allowing for training an effective Bias Amplifier.
-This is reflected by the final WGA ($90.03\%$).
-A similar behavior can be found for $\rho=0.800$ (WGA = $89.53\%$).
-As it was for Recipe I, $\rho=0.700$ results in a sensibly less effective Bias Amplifier due to the relevant percentage of conflicting samples among generated images.
-As a consequence, the final WGA is negatively affected ($73.52\%$).
-Input $\rho$ $0.950$ $0.900$ $0.800$ $0.700$
-Synthetic output $0.990$ $0.961$ $0.912$ $0.827$
-WGA $\bf 91.56$ $90.03$ $89.53$ $73.52\%$
-Table 5: Ablation on the number of training images for the Bias Amplifier, in terms of resulting WGA on Waterbirds, when using DDB’s Recipe II.
 Bias amplifier data requirements.
@@
 4.4, we want to ensure that applying Recipe II on an unbiased dataset does not degrade the final performance.
-Also Recipe II demonstrates to be applicable in unbiased scenarios, as it measures a final test accuracy of $83.87\%$ on the original test set of CIFAR10, which is slightly higher than the ERM baseline ($83.26\%$).
-Full Implementation Details of Recipe I (two-step)
+Also Recipe II demonstrates to be applicable in unbiased scenarios, as it measures a final test accuracy of $83.87\%$ on CIFAR10, which is slightly higher than the ERM baseline ($83.26\%$).
+Appendix B Full Implementation Details of Recipe I (two-step)
 For the hyperparameters of Recipe I, we use for each dataset the following configurations.
-Corrupted CIFAR-10: regardless of the specific $\rho$, we set batch size = 32, learning rate = $1\times e^{-4}$, weight decay = 0.01, and training is set to a total of 150 epochs.
 BAR: batch size = 32, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80.
@@
 The GDRO optimization is always run with the --robust and --reweight_groups flags.
-Full Implementation Details of Recipe II (end-to-end)
-For Corrupted CIFAR-10 and BAR, we use the same parameters reported in [6].
-Corrupted CIFAR-10: Regardless of the specific $\rho$, we set batch size = 256, learning rate = $5\times e^{-4}$, weight decay = $0.01$, and training is set to a total of 50 epochs.
-BAR: batch size = 128, learning rate = $1\times e^{-4}$, weight decay = 0.01, training epochs = 50.
+Appendix C Full Implementation Details of Recipe II (end-to-end)
+For BAR, we use the same parameters reported in [6]: batch size = 128, learning rate = $1\times e^{-4}$, weight decay = 0.01, training epochs = 50.
 Waterbirds: batch size = 128, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80.
 BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 0.01, training for 50 epochs.
-For the real-world datasets, we accumulate gradients for $8$ iterations ($16$ for BFFHQ) with mini-batches made of $16$ samples.
-Additional Biased Generations
-In this Section, we present $100$ additional per class generated images for each dataset utilized in the study, comprising three real-world and one synthetic dataset, as produced by our CDPM.
-With CIFAR10-C (Figure 6), the model proficiently synthesizes classes with specific corruptions, evidencing its ability to replicate synthetic bias patterns.
+For these recipes, we accumulate gradients for $8$ iterations ($16$ for BFFHQ) with mini-batches made of $16$ samples.
+Appendix D Additional Biased Generations
+In this Section, we present $100$ additional per class generated images for each dataset utilized in the study, as produced by our CDPM.
 In the Waterbirds dataset (Figure 6), the model adeptly captures and replicates the correlation between bird species and their background.
@@
 The results from the BAR dataset (Figure 8) exhibit the model’s adeptness in handling multiple bias patterns, preserving typical environmental contexts for different actions.
+Finally, the ImageNet-9 dataset (Figure 7) stereotypes the various classes and their backgrounds well.
 Overall, the model generates high-quality samples that are fidelity-biased across all datasets.
@@
 We leverage this characteristic to produce synthetic biased data aimed at enhancing the robustness of debiasing techniques.
-Figure 6: Synthetic bias-aligned image generations for each class of the CIFAR10-C dataset.
-Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
-Figure 7: Synthetic bias-aligned image generations for each class of the Waterbirds dataset.
-Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
-Figure 8: Synthetic bias-aligned image generations for each class of the BFFHQ dataset. Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
-Figure 9: Synthetic bias-aligned image generations for each class of the BAR dataset. Each grid shows 100 synthetic images for a class, revealing how the model aligns with dataset-specific biases.
-Bias Validation via Captions in Synthetic Images
+Figure 6: Synthetic bias-aligned image generations for each class of the Waterbirds dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
+Figure 7: Synthetic bias-aligned image generations for each class of the BAR dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
+Figure 8: Synthetic bias-aligned image generations for each class of the ImageNet-9 dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
+Figure 9: Synthetic bias-aligned image generations for each class of the BFFHQ dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
+Appendix E Bias Validation via Captions in Synthetic Images
 As an ulterior confirmation of bias in CDPM-generated images, besides two independent annotators, we implement an unsupervised analysis pipeline using the pre-trained BLIP-2 FlanT5XL model [40] for zero-shot image captioning without any task-specific prompting, for avoiding any guidance in the descriptions of the synthetic data.
@@
 Figure 11: Word frequency analysis from 1000 generated captions for BFFHQ classes ‘young’ (left) and ‘old’ (right), showing top 10 most frequent terms.
-Figure 12: Word frequency analysis from 1000 generated captions for BAR classes, showing top 10 most frequent terms for each class.
 (a) Class 0
@@
 (f) Class 5
+Figure 12: Word frequency analysis from 1000 generated captions for BAR classes, showing top 10 most frequent terms for each class.

### v4

@@
 Abstract
-Deep learning model effectiveness in classification tasks is often challenged by the quality and quantity of training data whenever they are affected by strong spurious correlations between specific attributes and target labels.
+The effectiveness of deep learning models in classification tasks is often challenged by the quality and quantity of training data whenever they are affected by strong spurious correlations between specific attributes and target labels.
 This results in a form of bias affecting training data, which typically leads to unrecoverable weak generalization in prediction.
-This paper aims at facing this problem by leveraging bias amplification with generated synthetic data: we introduce Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods of unsupervised model debiasing exploiting the inherent bias-learning tendency of diffusion models in data generation.
-Specifically, our approach adopts conditional diffusion models to generate synthetic bias-aligned images, which replace the original training set for learning an effective bias amplifier model that we subsequently incorporate into an end-to-end and a two-step unsupervised debiasing approach.
-By tackling the fundamental issue of bias-conflicting training samples memorization in learning auxiliary models, typical of this type of techniques, our proposed method beats current state-of-the-art in multiple benchmark datasets, demonstrating its potential as a versatile and effective tool for tackling bias in deep learning models.
-footnotetext: Correspondence: vito.paolo.pastore@unige.it
-Figure 1: Comparison of bias amplifier training approaches.
-(a) Traditional methods that use real biased datasets lead to the memorization of biased-conflicting samples, resulting in suboptimal auxiliary models.
-(b) Our proposed approach leverages conditional diffusion models to learn class-specific biases and amplify them into synthetic images.
-Such generations can replace the original training set for learning an effective Bias Amplifier, eliminating memorization effects by avoiding exposure to real data.
+This paper addresses this problem by leveraging bias amplification with generated synthetic data only: we introduce Diffusing DeBias (DDB), a novel approach acting as a plug-in for common methods of unsupervised model debiasing, exploiting the inherent bias-learning tendency of diffusion models in data generation.
+Specifically, our approach adopts conditional diffusion models to generate synthetic bias-aligned images, which fully replace the original training set for learning an effective bias amplifier model to be subsequently incorporated into an end-to-end and a two-step unsupervised debiasing approach.
+By tackling the fundamental issue of bias-conflicting training samples’ memorization in learning auxiliary models, typical of this type of technique, our proposed method outperforms the current state-of-the-art in multiple benchmark datasets, demonstrating its potential as a versatile and effective tool for tackling bias in deep learning models.
 1 Introduction
-Deep learning models have shown impressive results in image classification, yet their success depends heavily on training data quality and representativeness.
+Deep learning models have shown impressive results in various vision tasks, including image classification.
+However, their success is highly dependent on the quality and representativeness of the training data.
 When a deep neural network is trained on a biased dataset, so-called “shortcuts”, corresponding to spurious correlations (i.e., biases) between irrelevant attributes and labels, are learned instead of semantic class attributes, affecting model generalization and impacting test performances.
 In other words, the model becomes biased towards specific training sub-populations presenting biases, known as bias-aligned, while samples not affected by the bias are referred to as bias-conflicting or unbiased.
-In this scenario, several debiasing methods have been proposed [1, 2, 3, 4, 5], addressing the problem under different perspectives.
-Notably, the majority of available training samples are bias-aligned in a biased dataset (typically, $95\%$ or higher [6]).
-As such, common supervised debiasing strategies involve reweighting the training set to give more importance to the few available bias-conflicting data [5], for instance, by upsampling [7] or up-weighting [8] them.
-Conversely, unsupervised debiasing methods, which assume no prior availability of bias information, typically involve an auxiliary model to estimate and convey information on bias for reweighting the training set, employing their signal to guide the training of a debiased target model, while diminishing its tendency to learn bias shortcuts.
-To this end, there are several existing works [6, 1, 5, 9] exploiting intentionally biased auxiliary models, trained to overfit bias-aligned samples, thus being very confident in both correctly predicting bias-aligned training samples as well as misclassifying bias-conflicting ones [5].
-Such a distinct behavior between the two subpopulations is required to convey effective bias supervisory signals for target model debiasing, including gradients or per-sample loss values, which, e.g., should ideally be as high as possible for bias-conflicting and as low as possible for bias-aligned [6].
+In this scenario, where it is generally assumed that most of the training samples available in a biased data set are biased (typically, $95\%$ or higher nam2020learning ), several debiasing methods have been proposed addressing the problem from different points of view NEURIPS2022_75004615_LWBC ; tartaglione2023disentangling ; pastore2024lookingmodeldebiasinglens ; kim2021biaswap ; Lee_Park_Kim_Lee_Choi_Choo_2023 .
+When annotations on bias attributes are available, common supervised debiasing strategies involve reweighting the training set to give more importance to the few available bias-conflicting data Lee_Park_Kim_Lee_Choi_Choo_2023 , for example, by upsampling li2019repair or upweighting Sagawa*2020Distributionally them.
+Conversely, unsupervised debiasing methods, which assume no prior availability of bias information, typically involve an auxiliary model to estimate and convey information on bias for reweighting the training set, employing its signal to guide the training of a debiased target model, while diminishing its tendency to learn bias shortcuts.
+To this end, there are several existing works nam2020learning ; NEURIPS2022_75004615_LWBC ; Lee_Park_Kim_Lee_Choi_Choo_2023 ; sohoni2020no exploiting intentionally biased auxiliary models, trained to overfit bias-aligned samples, thus being very confident in both correctly predicting bias-aligned training samples as well as misclassifying bias-conflicting ones Lee_Park_Kim_Lee_Choi_Choo_2023 .
+Such a distinct behavior between the two subpopulations is required to convey effective bias supervisory signals for target model debiasing, including gradients or per-sample loss values, which, e.g., should ideally be as high as possible for bias-conflicting and as low as possible for bias-aligned nam2020learning .
 However, such desired behavior is severely hindered by bias-conflicting training samples, despite their low number.
-In fact, the low cardinality of this population impedes a suitable modeling of the correct data distribution, yet, they are sufficient to harm the estimation of a reliable bias-aligned data distribution, despite the latter having much higher cardinality.
-First, bias-conflicting samples in the training set act as noisy labels, affecting the bias learning for the auxiliary models [5].
-Furthermore, these models tend to overfit and memorize bias-conflicting training samples within very few epochs, failing to provide different learning signals for the two subpopulations [10].
-Despite available strategies, including the usage of the Generalized Cross Entropy (GCE) loss function [6, 3], ensembles of auxiliary models [1, 5], or even ad hoc solutions tailored to the specific dataset considered (e.g., training for only one epoch as in [11]), how to define a protocol for effective and robust auxiliary model training for unsupervised debiasing approaches remains an open issue.
-Remarkably, recent works [3, 5, 1] underline how the target model generalization is profoundly impacted by the protocol used to train such an auxiliary model.
-In principle, if one could remove all bias-conflicting samples from the training set, it would be possible to train a bias-capturing model robust to bias-conflicting samples’ memorization and interference, and capable of conveying ideal learning signals [5].
+Indeed, the low cardinality of this population impedes a suitable modeling of the correct data distribution, yet, they are sufficient to harm the estimation of a reliable bias-aligned data distribution, despite the latter having much higher cardinality.
+This scenario occurs under two main conditions.
+First, bias-conflicting samples in the training set act as noisy labels, affecting the bias learning for the auxiliary models Lee_Park_Kim_Lee_Choi_Choo_2023 .
+Second, these models tend to overfit and memorize bias-conflicting training samples within very few epochs, failing to provide different learning signals for the two subpopulations zarlenga2024efficient .
+Despite the available strategies, including the usage of the Generalized Cross Entropy (GCE) loss function nam2020learning ; pastore2024lookingmodeldebiasinglens , ensembles of auxiliary models NEURIPS2022_75004615_LWBC ; Lee_Park_Kim_Lee_Choi_Choo_2023 , or even ad hoc solutions tailored to the specific dataset considered (e.g., training for only one epoch or assuming the availability of a bias-annotated validation pmlr-v139-liu21f_JTT ), how to define a protocol for effective and robust auxiliary model training for unsupervised debiasing remains an open issue.
+Remarkably, recent works pastore2024lookingmodeldebiasinglens ; Lee_Park_Kim_Lee_Choi_Choo_2023 ; NEURIPS2022_75004615_LWBC underline how the target model generalization is profoundly impacted by the protocol used to train such an auxiliary model.
+In principle, if one could remove all bias-conflicting samples from the training set, it would be possible to train a bias-capturing model robust to bias-conflicting samples’ memorization and interference, and capable of conveying ideal learning signals Lee_Park_Kim_Lee_Choi_Choo_2023 .
 To this aim, our intuition is to elude these drawbacks by skipping the use of the actual training data in favor of another set of samples synthetically generated by an adequately trained diffusion model.
-Specifically, we propose to circumvent bias-conflicting interference on auxiliary models leveraging fully biased synthetic data exploiting intrinsic properties of Conditional Diffusion Probabilistic Models (CDPMs).
+Specifically, we propose to circumvent bias-conflicting interference on auxiliary models leveraging synthetic biased data obtained by exploiting Conditional Diffusion Probabilistic Models (CDPMs).
 We aim to learn a per-class biased image distribution, from which we sample an amplified synthetic bias-aligned subpopulation.
-The resulting synthetic bias-aligned samples are then exploited for training a Bias Amplifier model instead of the original actual training set, solving the issue of bias-conflicting overfitting and interference by construction, as the bias-capturing model does not see the original training set at all.
+The resulting synthetic bias-aligned samples are then exploited for training a Bias Amplifier model, instead of using the original actual training set, solving the issue of bias-conflicting overfitting and interference by construction, as the bias-capturing model does not see the original training set at all.
 We refer to the resulting approach as Diffusing DeBias (DDB).
-DDB’s intuition is indirectly supported by recent works studying the problem of biased datasets for image generation tasks  [12, 13], and showing how diffusion models’ generations can be biased, as a consequence of bias present in the training sets (See Sec.
+The intuition is indirectly supported by recent works studying the problem of biased datasets for image generation tasks d2024openbias ; kim2024training , and showing how diffusion models’ generations can be biased, as a consequence of bias present in the training sets (See Sec.
 2).
-This tendency can be seen as an undesired behavior of generative models, impacting the fairness of generated images, and recent papers have started to propose solutions for decreasing diffusion models’ inclination to learn biases [14, 13, 15].
-However, we claim that what is considered a problem for generative tasks can be turned into a feature for image classification model debiasing, as it allows the training of a robust auxiliary model that can ideally be plugged into various debiasing schemes, solving bias-conflicting training samples memorization, by construction (see Figure 1).
+This tendency can be seen as an undesirable behavior of generative models, which affects the fairness of generated images, and recent articles have started to propose solutions to decrease the inclination of diffusion models to learn biases DBLP:journals/corr/abs-2408-15094 ; kim2024training ; DBLP:conf/icb/PereraP23 .
+However, we claim that what is considered a drawback for generative tasks can be turned into an advantage for image classification’s unsupervised model debiasing, as it allows the training of a robust auxiliary model that can ideally be plugged into various debiasing schemes, solving bias-conflicting training samples memorization by construction.
 For this reason, we refer to DDB as a plug-in for debiasing methods in image classification.
-To prove its effectiveness and versatility, we incorporate DDB’s bias amplifier into two different debiasing frameworks, named Recipe I and Recipe II, able to: (i) extract subpopulations further used as pseudo-labels within the popular G-DRO [8] algorithm (Recipe I, Sec.
+To prove its effectiveness and versatility, we incorporate DDB’s bias amplifier into two different debiasing frameworks, which we refer to as Recipe I and Recipe II, including a two-step and an end-to-end approach (see Sec.
+2).
+Specifically, we exploit our bias-amplifier to: (i) extract subpopulations further used as pseudo-labels within the popular G-DRO Sagawa*2020Distributionally algorithm (Recipe I, Sec.
 3.3.1), and (ii) provide a loss function for the training set to be used within an end-to-end method (Recipe II, Sec.
 3.3.2).
-Both approaches outperform the state-of-the-art on popular benchmark datasets.
-To summarize, our main contributions can be summarized as follows:
-  • We introduce DDB, a novel and effective unsupervised debiasing framework that exploits intrinsic properties of CDPMs to learn per-class bias-aligned distributions in biased datasets, with unknown bias information.
-The resulting CDPM is used for generating synthetic bias-aligned images later employed for training a robust Bias Amplifier model, hence avoiding bias-conflicting training sample memorization and interference (Sec.
+Both approaches are competitive with the state-of-the-art on popular benchmark datasets.
+In summary, our main contributions can be regarded as follows:
+  • We introduce DDB, a novel and effective debiasing framework that exploits CDPMs to learn per-class bias-aligned distributions in biased datasets with unknown bias information.
+The resulting CDPM is used to generate synthetic bias-aligned images later used to train a robust bias amplifier model, thus avoiding real bias-conflicting training sample memorization and construction interference (Sec.
 3).
@@
 3.3.2) algorithm.
-  • Both proposed approaches prove to be effective, surpassing state-of-the-art results by a significant margin (Sec. 4) across four different benchmark datasets.
+Both proposed approaches prove to be effective, achieving state-of-the-art results (Sec.
+4) across popular benchmark datasets, characterized by different types of single or multiple biases.
 2 Related Work
-In this section, we provide a broad categorization of model debasing works.
+In this section, we provide a broad categorization of model debiasing works.
 Supervised Debiasing.
 Supervised methods exploit available prior knowledge about the bias (e.g., bias attribute annotations), indicating whether a sample exhibits a particular bias (or not).
-A common strategy consists of training a bias classifier to predict such attributes so that it can guide the target model to extract unbiased representations [16, 17].
-Alternatively, bias labels can be thought of as pre-defined encoded subgroups among target classes, allowing for the exploitation of robust training techniques such as G-DRO [8].
-At the same time, bias information can be employed to apply regularization schemes aiming at forcing invariance over bias features [18, 19].
+A common strategy consists of training a bias classifier to predict such attributes so that it can guide the target model to extract unbiased representations alvi2018turning ; Xie2017ControllableIT .
+Alternatively, bias labels can be thought of as pre-defined encoded subgroups among target classes, allowing for the exploitation of robust training techniques such as G-DRO Sagawa*2020Distributionally .
+At the same time, bias information can be employed to apply regularization schemes aiming at forcing invariance over bias features barbano2023unbiased ; tartaglione2021end .
 Unsupervised Debiasing.
@@
 Among such methods, a popular strategy consists in involving auxiliary models at any level of the debiasing process.
-We categorize these methods into two-step  [11, 9, 4]), and end-to-end  [6, 20]), in the following paragraphs.
+We categorize these methods into two-step  pmlr-v139-liu21f_JTT ; sohoni2020no ; kim2021biaswap , and end-to-end  nam2020learning ; li2022discover , in the following paragraphs.
 End-to-end approaches.
 Here, bias mitigation is performed online with dedicated optimization objectives.
-Employing techniques such as ensembling and noise-robust losses have been proposed, though they require careful tuning of all the hyperparameters, which may require the availability of validation sets with bias annotations to work.
+Employing techniques such as ensembling and noise-robust losses has been proposed, though they require careful tuning of all the hyperparameters, which may require the availability of validation sets with bias annotations to work.
 In Learning from Failure (LfF), Nam et al.
-[6] assume that bias features are learned earlier than task-related ones: they employ a bias-capturing model to focus on easier (bias-aligned) samples through the GCE loss [21].
+nam2020learning assume that bias features are learned earlier than task-related ones: they employ a bias-capturing model to focus on easier (bias-aligned) samples through the GCE loss zhang2018gce .
 At the same time, a debiased network is jointly trained to assign larger weights to samples that the bias-capturing model struggles to discriminate.
 In DebiAN, Li et al.
-[20] jointly train an auxiliary model (the discoverer) and a classifier (the target model) in an alternate scheme so that the discoverer can identify bias learned by the classifier during training, optimizing an ad-hoc loss function for bias mitigation.
+li2022discover jointly train an auxiliary model (the discoverer) and a classifier (the target model) in an alternate scheme so that the discoverer can identify bias learned by the classifier during training, optimizing an ad-hoc loss function for bias mitigation.
 Two-step methods.
-Here, the auxiliary model is employed in a first step to identify bias in training data, generally through pseudo-labeling, while the second stage consists in actual bias mitigation exploiting the inferred pseudo-labels.
-In [11], the auxiliary model is trained with standard ERM and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa: debiasing is brought on by up-sampling the predicted bias-conflicting samples.
+Here, the auxiliary model is employed in a first step to identify bias in training data, generally through pseudo-labeling, while the second stage consists of actual bias mitigation exploiting the inferred pseudo-labels.
+In pmlr-v139-liu21f_JTT , the auxiliary model is trained with standard ERM and then used in inference on the training set, considering misclassified samples as bias-conflicting and vice-versa: debiasing is brought on by up-sampling the predicted bias-conflicting samples.
 Notably, they must rely on a validation set with bias annotations to stop the auxiliary model’s training after a few epochs to avoid memorization and to tune the most effective upsampling factor.
-In [1], a set of auxiliary classifiers is ensembled (bootstrap ensembling) into a bias-commmittee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample.
-Finally, a recent trend involves the usage of vision-language models to identify bias, further exploiting bias predictions to mitigate biases in classification models [22, 23, 24].
-Among the described methods, a subset exploit the auxiliary model for bias amplification ([5, 6, 9]) under the general assumption that a bias amplification model would either misclassify or be less confident on bias-conflicting samples, with final debiasing performance strongly dependent on such behavior, and specifically, on how strongly the auxiliary model captures training bias [3, 5].
-However, this fundamental assumption only holds if bias-conflicting samples are not memorized during training, which is likely to happen in very few training epochs [10, 3], especially considering that the very few training bias-conflicting samples are more likely to be overfitted, as reported in [8, 25].
-One could argue that an annotated validation set could be used for properly regularizing the bias amplification model, but the latter cannot be assumed available a priori in real cases, this being en emerging argument of discussion among the community [10, 3].
-For this reason, in this work and differently from the previously cited research, we propose that syntethic bias-aligned samples could be used instead of the original training set, solving the bias-conflicting memorization issue by construction.
-However, an unsupervised real-world scenario poses a problem for generating a pure distribution of bias-aligned samples.
-As a solution, we propose to leverage diffusion models for generating a synthetic bias-aligned training set to be used for training a bias amplifier, inspired by recent works, showing how they can manifest distorted, unfair, or stereotypical representations deviating from the true underlying data distribution [26], when exposed to biases present in the training data  [15].
-While the vast majority of available works regarding generative models and debiasing refer to strategies for obtaining unbiased synthetic data generation [27, 28], up to our knowledge, we are the first to propose the usage of generative models for bias amplification.
-The closest available work, employing generative models for model debiasing, is Jung et al.
-[29], where the authors introduce a debiasing method exploiting a GAN [30] trained for style-transfer between several biased appearances, with a different objective.
+In NEURIPS2022_75004615_LWBC , a set of auxiliary classifiers is ensembled (bootstrap ensembling) into a bias-commmittee, proposing that debiasing can be performed using a weighted ERM, where the weights are proportional to the number of models in the ensemble misclassifying a certain sample.
+Finally, a recent trend involves the usage of vision-language models to identify bias, further exploiting bias predictions to mitigate biases in classification models eyuboglu2022domino ; kim2024discovering ; zhang2023diagnosing .
+Among the described methods, some exploit the auxiliary model for bias amplification (Lee_Park_Kim_Lee_Choi_Choo_2023 ; nam2020learning ; sohoni2020no ) under the general assumption that a bias amplification model would either misclassify or be less confident on bias-conflicting samples.
+In this case, the final debiasing performance is heavily dependent on such behavior, and specifically, on how strongly the auxiliary model captures the bias in training data pastore2024lookingmodeldebiasinglens ; Lee_Park_Kim_Lee_Choi_Choo_2023 .
+However, this fundamental assumption only holds if bias-conflicting samples are not memorized during training, which is likely to happen in very few training epochs zarlenga2024efficient ; pastore2024lookingmodeldebiasinglens , especially considering that the very few training bias-conflicting samples are more likely to be overfitted, as reported in Sagawa*2020Distributionally ; cao2019learning .
+One could argue that an annotated validation set could be used for properly regularizing the bias amplification model, but the latter cannot be assumed available a priori in real cases, being this an emerging argument of discussion among the community zarlenga2024efficient ; pastore2024lookingmodeldebiasinglens .
+For this reason, differently from the previously cited research, we propose in this work that synthetic bias-aligned samples could be used instead of the original training set, solving the bias-conflicting memorization issue by construction.
+However, a real-world scenario using data with unknown bias poses a problem for generating a pure distribution of bias-aligned samples.
+As a solution, we propose to leverage diffusion models for generating a synthetic bias-aligned training set to be used for training a bias amplifier.
+While the vast majority of available works regarding generative models and debiasing refer to strategies for obtaining unbiased synthetic data generation bhat2023debiasinggenerativemodelsusing ; gerych2023debiasing , to the best of our knowledge, we are the first to propose the use of generative models for obtaining an amplified biased distribution.
+The closest work employing generative models for model debiasing is Jung et al.
+jung2023fighting , where a debiasing method exploiting a GAN 10.1145/3422622 is introduced, which is trained for style-transfer between several biased appearances with different objectives.
 Here, a target model is debiased through contrastive learning between images affected by different biases for bias-invariant representations.
-In our work, we show that it is possible to leverage the intrinsic tendency of diffusion models to capture bias in data, using conditioned generations to infer the biased distribution of each target class, thus allowing the sampling of new synthetic images, sharing the same bias pattern of original train data.
-Our method provides a memorization-free bias amplification model, capable of providing precise supervisory signals for accurate model debiasing virtually compatible with any debiasing approach involving auxiliary models.
-Figure 2: Schematic representation of our DDB framework.
-The debiasing process consists of two key steps: (A) Diffusing the Bias uses a conditional diffusion model with classifier-free guidance to generate synthetic images that preserve training dataset biases, and (B) employs a Bias Amplifier firstly trained on such synthetic data, and subsequently used during inference to extract supervisory bias signals from real images.
-These signals are used to guide the training process of a target debiased model by designing two debiasing recipes (i.e., 2-step and end-to-end methods).
+Similar strategies, where bias-conflicting samples are augmented to support model debiasing, can be found in a few other works biasadv ; hwang2022selecmix ; dataaug1 .
+In dataaug1 , the problem of bias-conflicting memorization is stressed, proposing a bias-unsupervised strategy to augment and increase the diversity of bias-conflicting features, exploiting a bias amplifier analogously to nam2020learning for their identification, while improving the generalization of a target debiased model.
+In biasadv , an intentionally biased auxiliary model is exploited, which is trained on the biased dataset, supplementing bias-conflicting samples for the target debiased model.
+Here, images are generated by solving an optimization problem consisting of attacking the auxiliary model predictions while preserving the ones of the target debiased model.
+In hwang2022selecmix , bias-conflicting samples are again augmented exploiting the popular mix-up approach zhang2018mixupempiricalriskminimization .
+Specifically, an auxiliary model is employed to amplify easy-to-learn features, and mix-up pairs with either the same ground truth label, but different biased features (evaluated with clustering in the feature space of an intentionally biased model), or samples with different biased features (i.e., belonging to different clusters), but showing the same ground truth label.
+Differently from the previously described approaches, we are proposing an alternative strategy to avoid bias-conflicting memorization, while significantly improving the generalization of the target debiased model.
+In our work, we show how it is possible to leverage the intrinsic tendency of diffusion models to capture bias in data, using conditioned generations to infer the biased distribution of each target class, thus allowing the sampling of new synthetic images, sharing the same bias pattern of the original training data.
+Our method designs a memorization-free bias amplified model, capable of providing precise supervisory signals for accurate model debiasing, virtually compatible with any debiasing approach involving auxiliary models.
 3 The Approach
 Our proposed debiasing approach is schematically depicted in Figure 1.
-In this section, we provide at first a general description of the problem setting (Sec.
-3.1), and then, we illustrate in detail DDB’s two main components, which include bias diffusion (Sec.
+In this Section, we first describe the general problem setting (Sec.
+3.1), then detail DDB’s two main components: bias diffusion (Sec.
 3.2) and the two Recipes for model debiasing (Sec.
@@
 Let us consider a general data distribution $p_{\text{data}}$, typically encompassing multiple factors of variation and classes, and to build a dataset of images with the associated labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$ sampled from such a distribution.
-Let us also assume that the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors of variations, i.e.possible biases such as context, appearance, acquisition noise, viewpoint, etc..
+Let us also assume that the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors of variations, i.e., possible biases such as context, appearance, acquisition noise, viewpoint, etc..
 In this case, data will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors.
-This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not presenting the same biases.
+This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not present the same biases.
 In the same way, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$.
-Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or merely not annotated, in a realistic setting [13] In this context, biased and unbiased samples equivalently refer to bias-aligned and bias conflicting samples..
+Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or merely not annotated, in a realistic setting kim2024training .
 If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization.
-This is due to the strong correlation between $b$ and $y$, often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity [4, 8, 31]), which is dominating over the true target distribution semantics.
-It is important to notice that data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation [12].
-For instance, given a Conditional Diffusion Probabilistic Models (CDPM) modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ (with parameters $\phi$) that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we expect that its generations will be biased, as also stated in [12, 13].
-While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature [6]), a CDPM predominantly learns the biased distribution of a specific class, i.e., $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.
+This is due to the strong correlation between $b$ and $y$, often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity kim2021biaswap ; Sagawa*2020Distributionally ; nahon2023mining ), which is dominating over the true target distribution semantics.
+It is important to notice that data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation d2024openbias .
+For instance, given a Conditional Diffusion Probabilistic Models (CDPM) modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ (with parameters $\phi$) that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we expect that its generations will be biased, as also stated in d2024openbias ; kim2024training .
+While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature nam2020learning ), a CDPM predominantly learns the biased distribution of a specific class, i.e., $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.
 3.2 Diffusing the Bias
@@
 The Diffusion Process.
-The diffusion process progressively converts data into noise through a fixed Markov chain of $T$ steps [32].
+The diffusion process progressively converts data into noise through a fixed Markov chain of $T$ steps DBLP:conf/nips/HoJA20 .
 Given a data point $\mathbf{x}_{0}$, the forward process adds Gaussian noise according to a variance schedule $\{\beta_{t}\}_{t=1}^{T}$, resulting in noisy samples $\mathbf{x}_{1},\dots,\mathbf{x}_{T}$.
@@
 In cases where additional context or conditioning is available, such as a class label $y$, diffusion models can use this information to guide the reverse process, generating samples that better reflect the target attributes and semantics.
-Classifier-Free Guidance (CFG) [33] introduces a flexible conditioning approach, allowing the model to balance conditional and unconditional outputs without dedicated classifiers.
+Classifier-Free Guidance (CFG) DBLP:journals/corr/abs-2207-12598 introduces a flexible conditioning approach, allowing the model to balance conditional and unconditional outputs without dedicated classifiers.
 The CFG technique randomly omits conditioning during training (e.g., with probability $p_{\text{uncond}}=0.1$), enabling the model to learn both generation modalities.
@@
 In this work, we empirically show how CDPM learns and amplifies the underlying biased distribution when trained on a biased dataset with strong spurious correlations, allowing bias-aligned image generation.
+Figure 1: Schematic representation of our DDB framework.
+The debiasing process consists of two key steps: (A) Diffusing the Bias uses a conditional diffusion model with classifier-free guidance to generate synthetic images that preserve training dataset biases, and (B) employs a Bias Amplifier firstly trained on such synthetic data, and subsequently used during inference to extract supervisory bias signals from real images.
+These signals are used to guide the training process of a target debiased model by designing two debiasing recipes (i.e., 2-step and end-to-end methods).
 3.3 DDB: Bias Amplifier and Model Debiasing
@@
 This model can be exploited in either 2-step or end-to-end approaches, denoted here as Recipe I and Recipe II, respectively.
+(a) Overview of Recipe I’s 2-step debiasing approach.
+(b) Overview of Recipe II’s end-to-end debiasing approach.
+Figure 2: Comparison of two debiasing strategies: Recipe I and Recipe II.
 3.3.1 Recipe I: 2-step debiasing
-Figure 3: Overview of Recipe I’s 2-step debiasing approach.
-The adopted 2-step approach consists in 1) applying the auxiliary model trained on biased generated data to perform a bias pseudo-labeling, hence estimating bias-aligned/bias-conflict split of original actual data, and 2) apply a bias supervised method to train a debiased target model for classification.
-For the latter, we use the group DRO algorithm [8] (G-DRO) as a proven technique for the pure debiasing step.
+The adopted 2-step approach consists of 1) applying the auxiliary model trained on biased generated data to perform a bias pseudo-labeling, hence estimating bias-aligned/bias-conflict split of original actual data, and 2) applying a bias supervised method to train a debiased target model for classification.
+For the latter, we use the group DRO algorithm Sagawa*2020Distributionally (G-DRO) as a proven technique for the pure debiasing step.
 In other words, being in the unsupervised bias scenario where the real bias labels are unknown, we estimate bias pseudo-labels performing an inference step by feeding the trained BA with the original actual training data, and identifying as bias-aligned the correctly classified samples, and as bias-conflicting those misclassified.
-Among possible strategies to assign bias pseudo-labels, such as feature-clustering [9] or anomaly detection [3], we adopt a simple heuristic based on the BA misclassifications.
+Among possible strategies to assign bias pseudo-labels, such as feature-clustering sohoni2020no or anomaly detection pastore2024lookingmodeldebiasinglens , we adopt a simple heuristic based on the BA misclassifications.
 Specifically, given a sample $(\mathbf{x}_{i},y_{i},c_{i})$ with $c_{i}$ unknown pseudo-label indicating whether $\mathbf{x}_{i}$ is bias conflicting or aligned, we estimate bias-conflicting samples as
-$\hat{c}_{i}=\mathds{1}\left(\hat{y}_{i}\neq y_{i}~{}\land~{}\mathcal{L}(\hat{y}_{i},y_{i})~{}>~{}\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})\right)$ (3)
+$\hat{c}_{i}=\mathds{1}\left(\hat{y}_{i}\neq y_{i}~{}\land~{}\mathcal{L}(\hat{y}_{i},y_{i})~{}>~{}\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})\right),$ (3)
 where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA on the real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively, depending on the loss $\mathcal{L}$.
@@
 3).
-Once bias pseudo-labels over original training data are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 3.
+Once bias pseudo-labels over original training data are obtained, we plug in our estimate as group information for the G-DRO optimization, as schematically depicted in Figure 2(a).
 The above filtering operation refines the plain error set, restricting bias-conflicting sample selection to the hardest training samples, with potential benefits for the most difficult correlation settings ($\rho>0.99$).
@@
 3.3.2 Recipe II: end-to-end debiasing
-A typical end-to-end debiasing setting includes the joint training of the target debiasing model and one [6] or more [1, 5] auxiliary intentionally-biased models.
-Here, we design an end-to-end debiasing procedure, denoted as Recipe II, incorporating our BA by customizing a widespread general scheme, introduced in the Learning from Failure (LfF) method [6].
+A typical end-to-end debiasing setting includes the joint training of the target debiasing model and one nam2020learning or more NEURIPS2022_75004615_LWBC ; Lee_Park_Kim_Lee_Choi_Choo_2023 auxiliary intentionally-biased models.
+Here, we design an end-to-end debiasing procedure, denoted as Recipe II, incorporating our BA by customizing a widespread general scheme, introduced in the Learning from Failure (LfF) method nam2020learning .
 LfF leverages an intentionally-biased model trained using Generalized CE (GCE) loss to support the simultaneous training of a debiased model adopting the CE loss re-weighted by a per-sample relative difficulty score.
-Specifically, we replace the GCE biased model with our Bias Amplifier, which is frozen and only employed in inference to compute its loss function for each original training sample ($\mathcal{L}_{\text{bias\_amp}}$), as schematically represented in Figure 4.
-Such loss function is used to obtain a weighting factor for the target model loss function, defined as $r=\frac{\mathcal{L}_{\text{Bias\_Amp}}}{\mathcal{L}_{\text{debiasing}}+\mathcal{L}_{\text{Bias\_Amp}}}$.
+Specifically, we replace the GCE biased model with our Bias Amplifier, which is frozen and only employed in inference to compute its loss function for each original training sample ($\mathcal{L}_{\text{bias\_amp}}$), as schematically represented in Figure 2(b).
+Such a loss function is used to obtain a weighting factor for the target model loss function, defined as $r=\frac{\mathcal{L}_{\text{Bias\_Amp}}}{\mathcal{L}_{\text{debiasing}}+\mathcal{L}_{\text{Bias\_Amp}}}$.
 Coarsely speaking, $r$ should be low for bias-aligned and high for bias-conflicting samples.
-Figure 4: Overview of Recipe II’s end-to-end debiasing approach.
 4 Experiments
-In this Section, after the description of the benchmark datasets employed in our work, we provide general implementation and training details, followed by a discussion of the obtained results and a thorough ablation analysis.
-In the Supplementary Material, we report Recipe I and Recipe II full implementation details, more examples of generated images, and a quantitative analysis of the dataset bias captured from the CDPM with the assistance of a Vision-Language Model.
+In this Section, we report the description of the benchmark datasets employed, followed by a discussion of the obtained results and a thorough ablation analysis.
+In the Appendix, we report the general implementation (Sec.
+A.2) and training details, including the implementation details of Recipe I (Sec.
+A.3) and Recipe II (Sec.
+A.4), and more examples of the generated images (Sec.
+A.6).
 Figure 3: Examples of synthetic bias-aligned images across multiple datasets.
@@
 4.1 Datasets
-Our experiments utilize five distinct datasets: Waterbirds, Biased Flickr-Faces-HQ (BFFHQ), Biased Action Recognition (BAR), and ImageNet-9/ImageNet-A.
-Waterbirds  [8] is an image dataset exhibiting strong correlations ($\rho=0.950$) between bird species and background environments (e.g., water vs.
-land).
-It has 4,795 training images, 1,199 images for validation, and 5,794 for testing.
-BFFHQ [4] builds upon FFHQ [34] by introducing demographic biases for face recognition, while BAR  [6] is crafted to challenge action recognition models that associate specific actions with particular stereotypical contexts.
-For these two last datasets, we specifically refer to their original versions introduced respectively in [6] and [4]: BFFHQ has a total of 19,200 training images where  ( $y\in\{\text{young, old}\},~{}~{}b\in\{\text{female, male}\},~{}\rho=0.995$).
-Then, it provides 1,000 validation images (all bias-aligned) and 1,000 testing images with uniformly distributed labels and bias attributes.
-BAR presents 6 different classes, but does not provide explicit bias annotations or a validation set, with 1,941 images for training and 654 for testing.
-Finally, ImageNet-A consists of 7,500 real-world images from a 200-class subset of ImageNet, where deep learning models consistently struggle to make correct predictions.
-Originally introduced by Hendrycks et al.
-[35] to evaluate adversarial robustness, it is also employed as a model debiasing benchmark, usually as a bias-conflicting evaluation set for models trained on ImageNet-9 [36], a collection of images with 57,200 samples where categories are super-classes of ImageNet, put together to exhibit textural and shape biases [36].
-As such, both our CDPM and Bias Amplifier are trained on ImageNet-9.
-Then, we evaluate both recipes on ImageNet-A in terms of Average Accuracy, following the same protocol of [36, 1, 31]
-4.2 Implementation Details
-In our experiments, we generate $1000$ synthetic images for each target class using a CDPM implemented as a multi-scale UNet incorporating temporal and conditional embeddings for processing timestep information and class labels.
-For computational efficiency, input images are resized to $64\times 64$ pixels.
-Input images are normalized with respect to ImageNet’s mean and standard deviation.
-The CDPM is trained from scratch for $70,000$, $150,000$, $200,000$, $100.000$ iterations for the datasets BAR, Waterbirds, BFFHQ, and ImageNet-9 respectively.
-Batch size is set to $32$, and the diffusion process is executed over $1,000$ steps with a linear noise schedule (where $\beta_{1}=1e^{-4}$ and $\beta_{T}=0.028$).
-Optimization is performed using MSE loss and AdamW optimizer, with an initial learning rate of $1e^{-4}$, adjusted by a CosineAnnealingLR [37] scheduler with a warm-up period for the first $10\%$ of the total iterations.
-As per the Bias Amplifier training, we use synthetic images from CDPM.
-A Densenet-121 [38] with ImageNet pre-training is employed across all datasets, featuring a single-layer linear classification head.
-The training uses regular Cross-Entropy loss function and the AdamW optimizer [39].
-For BFFHQ and BAR, the learning rate is $1e^{-4}$, and $5e^{-4}$ for Waterbirds, and ImageNet-9.
-Training spans 50 epochs with weight decay $\lambda=0.01$ except for BFFHQ, which uses $\lambda=1.0$ for 100 epochs.
-We apply standard basic data augmentation strategies (following [4, 6]), including random crops and horizontal flips.
-For our debiasing recipes, we rely on the same backbones used in the existing literature, i.e.ResNet-18 for BFFHQ, BAR, ImageNet-9/A, and a ResNet-50 for Waterbirds [6, 4, 23].
-In Recipe II, $\gamma$ (Equation 3) is set to 3 in all our main experiments.
-Finally, similarly to [20, 6, 3], we do not rely on a validation set for regularization.
-This choice is related to real-world application suitability, where no guarantees exist regarding the degree of spurious correlations present in a held-out subset of data used as validation, possibly resulting in detrimental regularization [10].
+Our experiments utilize six distinct datasets: three typical benchmark datasets, including Waterbirds Sagawa*2020Distributionally , Biased Flickr-Faces-HQ (BFFHQ) kim2021biaswap , and Biased Action Recognition (BAR), taken from the original version as introduced in nam2020learning .
+On top of these benchmark datasets, we are interested in evaluating our approach in challenging scenarios involving real-world images with natural biases and datasets where multiple shortcuts are present.
+For this reason, we include ImageNet-9 bahng2019rebias /ImageNet-A hendrycks2021natural , and UrbanCars li2023whac in our experiments.
+More dataset details are in Sec.
+A.1.
 4.2 Results
@@
 First, we present qualitative image generation results across multiple datasets (see Fig.
-5), to validate the effectiveness of our bias diffusion approach.
-The generated samples maintain good fidelity and capture the typical bias patterns of each training set per class, confirming our conditioned diffusion process’s ability to learn and diffuse inherent biases.
-This is confirmed by screening the generated images by 2 independent annotators, who were asked to verify (or deny) the presence of the bias in each synthetic sample.
-In addition, we also carried out an objective evaluation of the generated image content by adopting an image captioner, specifically BLIP-2 [40].
+3) to validate the effectiveness of our bias diffusion approach.
+The generated samples maintain good fidelity and capture the typical bias patterns of each training set per class, confirming the conditioned diffusion process’s ability to learn and diffuse inherent biases.
+This is confirmed by screening the generated images by three independent annotators, who were asked to verify (or deny) the presence of the bias in each synthetic sample, where possible.
+Specifically, for Waterbirds, the reported average bias-aligned proportion of synthetic generated images is $98.35\%$, in BAR it is $99.20\%$, and $99.70\%$ is obtained for BFFHQ.
+All these values are higher than the known proportions of bias in the original datasets.
+In addition, focusing on Waterbirds, we also evaluated the generated image content by adopting an image captioner, specifically BLIP-2 DBLP:conf/icml/0008LSH23 .
 We estimate a caption for each image per class and count the frequency of every word.
-Suppl.
-Material (Sec.
-A.E) shows a histogram of the detected keywords related to the Waterbirds dataset, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced.
-The same behavior also emerges in other datasets used in this study, where a consistent description of the bias attributes can be noted, as reported in Sec.
-A.E.
-These results highlight our CDPM’s capability to generate synthetic bias-aligned data, setting the stage for our subsequent debiasing techniques.
+Supplementary Material (Sec.
+A.7) shows a histogram of the detected keywords, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced.
 4.2.2 Classification accuracy
 Table 1 reports the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets.
-We report Worst Group Accuracy (WGA) for Waterbirds, as in [8, 5].
-For BFFHQ, we report the average and the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following [4, 20].
+We report Worst Group Accuracy (WGA) for Waterbirds, as in Sagawa*2020Distributionally ; Lee_Park_Kim_Lee_Choi_Choo_2023 .
+For BFFHQ, we report the average and the conflicting accuracy (i.e., the average accuracy computed only on the bias-conflicting test samples), following kim2021biaswap ; li2022discover .
 BAR and ImageNet-A do not provide bias-attribute annotations, hence, we report just the average test accuracy.
 For simplicity, throughout this section, we will refer to Recipe I and Recipe II as DDB-I and DDB-II, respectively.
-It’s worth noticing that, for BAR and BFFHQ, we report benchmark results related to works that employ the same versions as ours.
+It is worth noticing that, for BAR and BFFHQ, we report benchmark results related to works that employ the same versions as ours.
 Method Unsup Waterbirds BAR BFFHQ ImageNet-A
@@
 […]
-LISA [41] $89.20$
+LISA yao2022improving $89.20$
 […]
@@
 […]
-Park et al. [44] $71.68$
+Park et al. park2024enhancing $71.68$
 […]
-[…]
-CDvG+LfF [29] $84.80$ $62.20\text{\scriptsize{$\pm$0.45}}$ $34.60$
+CDvG+LfF jung2023fighting $84.80$ $62.20\text{\scriptsize{$\pm$0.45}}$ $34.60$
 […]
@@
 […]
-DDB-I
-(w/ plain err. set)
 […]
-Table 1: Results of DDB Recipes I and II on benchmark datasets.
-Ticks () indicate that the method is bias-unsupervised.
-We highlight in bold the best results from unsupervised debiasing works, second-best are underlined.
+Table 1: Results of DDB Recipes I and II on benchmark datasets. Ticks () mark bias-unsupervised methods. Best unsupervised results are in bold; second-best are underlined.
 Across all the considered benchmark datasets, both DDB recipes show state-of-the-art results (see Table 1).
-Notably, in Waterbirds, DDB’s WGA is higher than G-DRO supervised (91.56% vs.
-91.40% of G-DRO [8]) and all the other unsupervised methods’ ones.
-DDB-I is providing slightly worse performance with respect to DDB-II, but still better than the state of the art.
-In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN [20]), in terms of conflicting accuracy.
-In BFFHQ, our DDB-I reaches an accuracy of 74.67%, with an average improvement of 1% when compared to ETF-Debias [43] (73.60%), the best state-of-the-art results to date, and provides a boost of 7.45% over the recent DeNetDM [45].
+Notably, in Waterbirds, DDB’s WGA is higher than all the other state-of-the-art unsupervised methods’ accuracies, as well as of that of the supervised G-DRO Sagawa*2020Distributionally (91.56% vs.
+91.40% of G-DRO).
+DDB-I is providing slightly worse performance than DDB-II, but still better than the state of the art.
+In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN li2022discover ), in terms of conflicting accuracy.
+In BFFHQ, our DDB-I reaches an accuracy of 74.67%, with an average improvement of 1% when compared to the state-of-the-art ETF-Debias wang2024navigate (73.60%).
+Method ID.ACC $\uparrow$ BG-Gap $\downarrow$ CoObj-Gap $\downarrow$ BG-CoObj-Gap $\downarrow$
+GroupDRO Sagawa*2020Distributionally $91.60$ $10.90$ $3.60$ $16.40$
+LLE li2023whac $96.70$ $2.10$ $2.70$ $5.90$
+LfF nam2020learning $97.20$ $11.60$ $18.40$ $63.20$
+JTT pmlr-v139-liu21f_JTT $95.90$ $8.10$ $13.30$ $40.10$
+EIIL creager2021environment $95.50$ $4.20$ $24.70$ $44.90$
+DebiAN li2022discover $\underline{98.00}$ $14.90$ $10.50$ $69.00$
+[…]
+[…]
+Table 2: Comparison of deep learning methods on the UrbanCars biased dataset.
+ID.ACC is in-distribution accuracy (the higher, the better); other three metrics show the accuracy gap between each minority group and ID.ACC (the lower, the better).
+For UrbanCars (see Table 2), presenting multiple biases, we adopted the same metrics of li2023whac (see Suppl.
+Material, Sec.
+A.1).
+DDB-I shows overall the most balanced performance across bias-conflicting samples, reporting an improvement over LLE li2023whac of $0.25\%$ for BG-Gap, $2.18\%$ on CoObj-Gap, and $5.78\%$ for BG-CoObj-Gap.
+DDB-II provides a significant improvement over the original LfF, but struggles more with the background-conflicting samples than DDB-I.
 4.3 Ablation analysis
@@
 Without losing generality, all ablations are performed on Waterbirds and exploit Recipe I for model debiasing.
-Additional ablations on Recipe II can be found in the Suppl.
-Material.
+Additional ablations on Recipe II can be found in the Supp.
+Mat.
 The reported analyses regard several aspects of the data generation process with respect to the Bias Amplifier, the effects of the filtering option in Recipe I, and the behavior of DDB on unbiased data.
 Bias Amplifier data requirements.
-All our main results are obtained by training a BA on $1000$ synthetic images per class.
-Here, we want to measure how this choice impacts the effectiveness of our Bias Amplifier.
+All our main results are obtained by training a BA on $1000$ synthetic images per class, and we want here to measure how this choice impacts its effectiveness.
 From Table 3 (left), we observe that even 100 synthetic images are enough for good debiasing performance despite not being competitive with top accuracy scores.
-As we increase the number of training images, WGA improves, reaching the best performance for $2000$ images per class.
+Increasing the number of training images improves WGA, reaching the best performance for $1000$ images per class.
 CFG strength and generation bias.
@@
 Table 3 (right) shows a limited impact on debiasing performance.
-WGA variations along different $w$ values are limited, with the best performance obtained for $w=1$.
+WGA variations along different $w$ values equal or larger than 1 are limited, with the best performance obtained for $w=1$.
 # Synth. images for BA training
@@
 Our results show a trade-off between the two alternatives.
-The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings ($\rho=0.950$) (e.g., Waterbirds), gaining an advantage from such a coarser selection.
-On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios, like BFFHQ ($\rho=0.995$).
-Bias-Identification Accuracy To frame the goodness of our Bias Amplifier in correctly identifying aligned and conflicting training samples, we provide the accuracy of such process when considering samples misclassified by the BA.
-Specifically, we report a bias-identification accuracy of $89.00\%$ for Waterbirds and $96.00\%$ on BFFHQ.
-Our very high accuracy supports the effectiveness of using the proposed protocol, which is not dependent from any careful regularization, bias-annotated validation sets, or large ensembles of auxiliary classifiers.
+The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings (e.g., like in Waterbirds, $\rho=0.950$), gaining an advantage from such a coarser selection.
+On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios, like in BFFHQ ($\rho=0.995$).
+Bias-Identification Accuracy.
+To frame the goodness of our BA in correctly identifying aligned and conflicting training samples, we provide the accuracy of such a process when considering samples misclassified by it.
+Specifically, we reach a high bias-identification accuracy of $89.00\%$ for Waterbirds and $96.00\%$ on BFFHQ, supporting the effectiveness of the proposed protocol, which is not dependent on any careful regularization, bias-annotated validation sets, or large ensembles of auxiliary classifiers.
 DDB impact on unbiased dataset.
 In real-world settings, it is generally unknown whether bias is present in datasets or not.
-Consequently, an effective approach, while mitigating bias dependency in the case of a biased dataset should also not degrade the performance when the bias is not present.
-This ablation study aims to assess DDB from this perspective.
-To obtain this insight, we apply the entire DDB pipeline (adopting Recipe I) on a common image dataset such as CIFAR-10  [46], comparing it with traditional ERM training with CE loss.
+Consequently, an effective approach, while mitigating bias dependency in the case of a biased dataset, should also not degrade the performance when the bias is not present.
+To evaluate DDB in this regard, we apply the entire pipeline (adopting Recipe I) on a common image dataset such as CIFAR-10  krizhevsky2009learning , comparing it with traditional ERM training with CE loss.
 In this controlled case, Recipe I provides a test accuracy of $84.16\%$, with a slight improvement of $\sim 1\%$ compared to traditional ERM ($83.26\%$) employing the same target model (ResNet18).
-This result shows that our Bias Amplifier can still provide beneficial signals to the target model in standard learning scenarios, potentially favoring the learning of the most challenging samples.
+This result shows that our Bias Amplifier can still provide beneficial signals to the target model in standard scenarios, potentially favoring the learning of the most complex samples.
 5 Conclusions
 In this work, we present Diffusing DeBias (DDB), a novel debiasing framework capable of learning the per-class bias-aligned training data distribution, leveraging a CDPM.
-Synthetic images sampled from the inferred biased distribution are used to train a Bias Amplifier, employed to provide a supervisory signal for training a debiased target model.
-The usage of synthetic bias-aligned data permits avoidance of detrimental effects typically impacting auxiliary models’ training in unsupervised debiasing schemes, such as bias-conflicting sample overfitting and interference, which lead to suboptimal generalization performance of the debiased target model.
-DDB is effective and versatile, acting as a plug-in for unsupervised debiasing methods.
-In this work, we design a two-step and an end-to-end debiasing recipes incorporating DDB’s Bias Amplifier: they outperform state-of-the-art works by a margin on four popular biased benchmark datasets, including the most recent relying on vision-language models [23, 47].
-Finally, while effectively mitigating bias dependency for biased datasets, DDB does not degrade performance when used with unbiased data, making it suitable for debiasing in real-world applications.
+Synthetic images sampled from the inferred biased distribution are used to train a Bias Amplifier, employed to provide a supervisory signal for training a debiased model.
+The usage of synthetic bias-aligned data prevents the detrimental effects typically impacting auxiliary models’ training in unsupervised debiasing schemes, such as bias-conflicting sample overfitting and interference, which lead to suboptimal generalization performance of the debiased target model.
+DDB results effective and versatile, acting as a plug-in for unsupervised debiasing methods.
+In this work, we incorporate DDB’s Bias Amplifier in a two-step and an end-to-end debiasing recipes.
+leading to accuracies outperforming state-of-the-art works by a margin on popular biased benchmark datasets (considering single and multiple biases), including recent works relying on vision-language models kim2024discovering ; ciranni2024say .
+Finally, while effectively mitigating bias dependency for biased datasets, DDB does not degrade performance when used with unbiased data, making it suitable to be adopted as an effective general-purpose standard training procedure.
 Limitations.
@@
 Another limitation inherited from diffusion models is that the generation quality may decrease if the training dataset is too small, impacting DDB applicability for very small-scale datasets.
-Appendix
-Appendix A Ablation Studies with DDB’s Recipe II
+Appendix A Technical Appendices and Supplementary Material
+A.1 Datasets
+  • Waterbirds  Sagawa*2020Distributionally is an image dataset exhibiting strong correlations ($\rho=0.950$) between bird species and background environments (e.g., water vs.
+land).
+It has 4,795 training images, 1,199 images for validation, and 5,794 for testing.
+[…]
+  • We utilize the original BAR version as introduced in nam2020learning .
+The dataset presents 6 different classes, but does not provide explicit bias annotations or a validation set, with 1,941 images for training and 654 for testing.
+  • ImageNet-A consists of 7,500 real-world images from a 200-class subset of ImageNet, where deep learning models consistently struggle to make correct predictions.
+Originally introduced by Hendrycks et al.
+hendrycks2021natural to evaluate adversarial robustness, it is also employed as a model debiasing benchmark, usually as a bias-conflicting evaluation set for models trained on ImageNet-9 bahng2019rebias , a collection of images with 57,200 samples where categories are super-classes of ImageNet, put together to exhibit textural and shape biases bahng2019rebias .
+As such, both our CDPM and Bias Amplifier are trained on ImageNet-9.
+Then, we evaluate both recipes on ImageNet-A in terms of Average Accuracy, following the same protocol of bahng2019rebias ; NEURIPS2022_75004615_LWBC ; nahon2023mining .
+  • UrbanCars has been released in li2023whac and contains 8000 training images with the target represented by urban and country cars.
+The dataset is built starting from the original $196$ classes of Stanford Cars stanfordcars , while synthetically incorporating two different shortcuts, background, from the Places dataset zhou2017places , and Co-Occurring Objects (CoObj), from the LVIS one gupta2019lvis .
+In UrbanCars, $90.25\%$ of samples present both background and CoObj attributes aligned with the target class biases, $9.5\%$ of training samples are conflicting with respect to only one of the two attributes, while $0.25\%$ are conflicting to both.
+The evaluation metrics employed for this datasets are the one introduced in li2023whac , namely ID.ACC, BG-Gap, CoObj-Gap, BGCoObj-Gap.
+ID.ACC is the in-distribution accuracy (the weighted per-class accuracy using the training set per-group subpopulation proportions).
+The other metrics represent the test accuracy gap of each minority subgroup with respect to ID.ACC.
+A.2 Implementation Details
+In our experiments, we generate $1000$ synthetic images for each target class using a CDPM implemented as a multi-scale UNet incorporating temporal and conditional embeddings for processing timestep information and class labels.
+For computational efficiency, input images are resized to $64\times 64$ pixels.
+Input images are normalized with respect to ImageNet’s mean and standard deviation.
+The CDPM is trained from scratch for $70,000$, $150,000$, $200,000$, $100.000$, $100.000$ iterations for the datasets BAR, Waterbirds, BFFHQ, ImageNet-9, and UrbanCars respectively.
+Batch size is set to $32$, and the diffusion process is executed over $1,000$ steps with a linear noise schedule (where $\beta_{1}=1e^{-4}$ and $\beta_{T}=0.028$).
+Optimization is performed using MSE loss and AdamW optimizer, with an initial learning rate of $1e^{-4}$, adjusted by a CosineAnnealingLR DBLP:conf/iclr/LoshchilovH17 scheduler with a warm-up period for the first $10\%$ of the total iterations.
+As per the Bias Amplifier training, we use synthetic images from CDPM.
+A Densenet-121 huang2017densely with ImageNet pre-training is employed across all datasets, featuring a single-layer linear classification head.
+The training uses regular Cross-Entropy loss function and the AdamW optimizer loshchilov2018decoupled .
+For BFFHQ, BAR, and UrbanCars, the learning rate is $1e^{-4}$, and $5e^{-4}$ for Waterbirds, and ImageNet-9.
+Training spans 50 epochs with weight decay $\lambda=0.01$ except for BFFHQ, which uses $\lambda=1.0$ for 100 epochs.
+We apply standard basic data augmentation strategies (following kim2021biaswap ; nam2020learning ), including random crops and horizontal flips.
+For our debiasing recipes, we rely on the same backbones used in the existing literature, i.e.ResNet-18 for BFFHQ, BAR, ImageNet-9/A, and a ResNet-50 for Waterbirds and UrbanCars nam2020learning ; kim2021biaswap ; kim2024discovering ; li2023whac .
+In Recipe II, $\gamma$ (Equation 3) is set to 3 in all our main experiments.
+Finally, similarly to li2022discover ; nam2020learning ; pastore2024lookingmodeldebiasinglens , we do not rely on a validation set for regularization.
+This choice is related to real-world application suitability, where no guarantees exist regarding the degree of spurious correlations present in a held-out subset of data used as validation, possibly resulting in detrimental regularization zarlenga2024efficient .
+A.3 Full Implementation Details of Recipe I (two-step)
+For the hyperparameters of Recipe I, we use for each dataset the following configurations.
+BAR: batch size = 32, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80.
+Waterbirds: batch size = 128, learning rate = $5\times e^{-4}$, weight decay = 1.0, training for 60 epochs.
+BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 1.0, training for 60 epochs.
+The GDRO optimization is always run with the --robust and --reweight_groups flags.
+For UrbanCars, presenting two biases, we apply Recipe I as it is, considering the filtered correct and wrong classifications in each class as 4 subgroups during GDRO optimization.
+We then consider all the available 8 subgroups at evaluation time.
+A.4 Full Implementation Details of Recipe II (end-to-end)
+For BAR, we use the same parameters reported in nam2020learning : batch size = 128, learning rate = $1\times e^{-4}$, weight decay = 0.01, training epochs = 50.
+Waterbirds: batch size = 128, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80.
+BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 0.01, training for 50 epochs.
+For these recipes, we accumulate gradients for $8$ iterations ($16$ for BFFHQ) with mini-batches made of $16$ samples.
+A.5 Ablation Studies with DDB’s Recipe II
 This appendix section focuses on additional ablation studies, dedicated to our Recipe II.
@@
 Similarly to what is shown in Sec.
-4.4, we want to ensure that applying Recipe II on an unbiased dataset does not degrade the final performance.
+4.3, we want to ensure that applying Recipe II on an unbiased dataset does not degrade the final performance.
 Also Recipe II demonstrates to be applicable in unbiased scenarios, as it measures a final test accuracy of $83.87\%$ on CIFAR10, which is slightly higher than the ERM baseline ($83.26\%$).
-Appendix B Full Implementation Details of Recipe I (two-step)
-For the hyperparameters of Recipe I, we use for each dataset the following configurations.
-BAR: batch size = 32, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80.
-Waterbirds: batch size = 128, learning rate = $5\times e^{-4}$, weight decay = 1.0, training for 60 epochs.
-BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 1.0, training for 60 epochs.
-The GDRO optimization is always run with the --robust and --reweight_groups flags.
-Appendix C Full Implementation Details of Recipe II (end-to-end)
-For BAR, we use the same parameters reported in [6]: batch size = 128, learning rate = $1\times e^{-4}$, weight decay = 0.01, training epochs = 50.
-Waterbirds: batch size = 128, learning rate = $5\times e^{-5}$, weight decay = 0.01, training epochs = 80.
-BFFHQ: batch size = 256, learning rate = $5\times e^{-5}$, weight decay = 0.01, training for 50 epochs.
-For these recipes, we accumulate gradients for $8$ iterations ($16$ for BFFHQ) with mini-batches made of $16$ samples.
-Appendix D Additional Biased Generations
+A.6 Additional Biased Generations
 In this Section, we present $100$ additional per class generated images for each dataset utilized in the study, as produced by our CDPM.
@@
 Figure 4: Synthetic bias-aligned image generations for each class of the Waterbirds dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
+Figure 5: Synthetic bias-aligned image generations for each class of the UrbanCars dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
 Figure 6: Synthetic bias-aligned image generations for each class of the BAR dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
@@
 Figure 8: Synthetic bias-aligned image generations for each class of the BFFHQ dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
-Appendix E Bias Validation via Captions in Synthetic Images
-As an ulterior confirmation of bias in CDPM-generated images, besides two independent annotators, we implement an unsupervised analysis pipeline using the pre-trained BLIP-2 FlanT5XL model [40] for zero-shot image captioning without any task-specific prompting, for avoiding any guidance in the descriptions of the synthetic data.
+A.7 Bias Validation via Captions in Synthetic Images
+As an ulterior confirmation of bias in CDPM-generated images, besides two independent annotators, we implement an unsupervised analysis pipeline using the pre-trained BLIP-2 FlanT5XL model DBLP:conf/icml/0008LSH23 for zero-shot image captioning without any task-specific prompting, for avoiding any guidance in the descriptions of the synthetic data.
 Our method processes the resulting captions through stop word removal and frequency analysis to reveal underlying biases without relying on predetermined categories or supervised classification models.
 Figure 9 presents the keyword frequency distribution for the Waterbirds dataset, where the most prevalent terms naturally correspond to known bias attributes (e.g., environmental contexts).
-The same is true for the keyword frequency distribution of the BFFHQ dataset (Figure 11) and BAR (Figure 12) where the most prevalent terms are related to the gender and the context respectively.
 These findings confirm that generated images are bias-aligned.
@@
 Figure 9: Word frequency analysis from 1000 generated captions for Waterbirds classes ‘landbird’ (left) and ‘waterbird’ (right), showing top 10 most frequent terms.
-Figure 11: Word frequency analysis from 1000 generated captions for BFFHQ classes ‘young’ (left) and ‘old’ (right), showing top 10 most frequent terms.
-(a) Class 0
-(b) Class 1
-(c) Class 2
-(d) Class 3
-(e) Class 4
-(f) Class 5
-Figure 12: Word frequency analysis from 1000 generated captions for BAR classes, showing top 10 most frequent terms for each class.

### v5

@@
 By tackling the fundamental issue of bias-conflicting training samples’ memorization in learning auxiliary models, typical of this type of technique, our proposed method outperforms the current state-of-the-art in multiple benchmark datasets, demonstrating its potential as a versatile and effective tool for tackling bias in deep learning models.
+Code is available at https://github.com/Malga-Vision/DiffusingDeBias.
 1 Introduction
@@
 Specifically, an auxiliary model is employed to amplify easy-to-learn features, and mix-up pairs with either the same ground truth label, but different biased features (evaluated with clustering in the feature space of an intentionally biased model), or samples with different biased features (i.e., belonging to different clusters), but showing the same ground truth label.
+Other works have also explored generative models for debiasing.
+For instance, BiaSwap kim2021biaswap utilizes image translation techniques to swap bias attributes between images, aiming to create more balanced training data.
+However, methods that attempt to create synthetic bias-conflicting samples kim2021biaswap ; biasadv may require high semantic fidelity, while our method creates a functionally biased dataset, which we believe to be a simpler and more robust objective.
 Differently from the previously described approaches, we are proposing an alternative strategy to avoid bias-conflicting memorization, while significantly improving the generalization of the target debiased model.
@@
 3.1 Problem Setting
-Let us consider a general data distribution $p_{\text{data}}$, typically encompassing multiple factors of variation and classes, and to build a dataset of images with the associated labels ${}{\mathcal{D}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}}$ sampled from such a distribution.
-Let us also assume that the sampling process to obtain $\mathcal{D}$ is not uniform across latent factors of variations, i.e., possible biases such as context, appearance, acquisition noise, viewpoint, etc..
-In this case, data will not faithfully capture the true data distribution ($p_{\text{data}}$) just because of these bias factors.
-This phenomenon deeply affects the generalization capabilities of deep neural networks in classifying unseen examples not present the same biases.
-In the same way, we could consider $\mathcal{D}$ as the union between two sets, i.e.$\mathcal{D}=\mathcal{D}_{\text{unbiased}}\bigcup\mathcal{D}_{\text{biased}}$.
-Here, the elements of $\mathcal{D}_{\text{unbiased}}$ are uniformly sampled from $p_{\text{data}}$ and, in $\mathcal{D}_{\text{biased}}$, they are instead sampled from a conditional distribution $p_{\text{data}}\left(\mathbf{x},y\>|\>b\right)$, with $b\in B$ being some latent factor (bias attribute) from a set of possible attributes $B$, likely to be unknown or merely not annotated, in a realistic setting kim2024training .
-If $|\mathcal{D}_{\text{biased}}|\gg|\mathcal{D}_{\text{unbiased}}|$, optimizing a classification model $f_{\theta}$ over $\mathcal{D}$ likely results in biased predictions and poor generalization.
-This is due to the strong correlation between $b$ and $y$, often called spurious correlation, and denoted as $\rho(y,b)$, or just $\rho$ for brevity kim2021biaswap ; Sagawa*2020Distributionally ; nahon2023mining ), which is dominating over the true target distribution semantics.
-It is important to notice that data bias is a general problem, not only affecting classification tasks but also impacting several others such as data generation d2024openbias .
-For instance, given a Conditional Diffusion Probabilistic Models (CDPM) modeled as a neural network $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)$ (with parameters $\phi$) that learns to approximate a conditional distribution $p\left(\mathbf{x}\>|\>y\right)$ from $\mathcal{D}$, we expect that its generations will be biased, as also stated in d2024openbias ; kim2024training .
-While this is a strong downside for image-generation purposes, in this work, we claim that when $\rho(y,b)$ is very high (e.g.$\geq 0.95$, as generally assumed in model debiasing literature nam2020learning ), a CDPM predominantly learns the biased distribution of a specific class, i.e., $\tilde{g}_{\phi}(\mathbf{x}~{}|~{}y)\approx p\left(\mathbf{x}\>|\>b\right)$ rather than $p\left(\mathbf{x}\>|\>y\right)$.
+[…]
+[…]
 3.2 Diffusing the Bias
@@
 This model can be exploited in either 2-step or end-to-end approaches, denoted here as Recipe I and Recipe II, respectively.
+This section describes the two debiasing recipes we employ in this work to demonstrate the effectiveness of our BA as a plug-in for unsupervised model debiasing.
 (a) Overview of Recipe I’s 2-step debiasing approach.
@@
 where $\mathds{1}$ is the indicator function, $\mathcal{L}$ is the CE loss of the BA on the real sample, and $\mu_{n}$ and $\sigma_{n}$ represent the average training loss and its standard deviation, respectively, depending on the loss $\mathcal{L}$.
-Together with the multiplier $\gamma~{}\in\mathbb{N}$, this condition defines a sort of filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss increased by a quantity corresponding to a certain z-score of the per-sample training loss distribution (${}{\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})}$ in Eq.
+Together with the multiplier $\gamma$$\in\mathbb{N}$, this condition defines a sort of filter over misclassified samples, considering them as conflicting only if their loss is also higher than the mean loss increased by a quantity corresponding to a certain z-score of the per-sample training loss distribution ($\mu_{n}(\mathcal{L})+\gamma\sigma_{n}(\mathcal{L})$} in Eq.
 3).
@@
 A.3) and Recipe II (Sec.
-A.4), and more examples of the generated images (Sec.
-A.6).
-Figure 3: Examples of synthetic bias-aligned images across multiple datasets.
-Each grid shows synthetic images for specific classes across biased datasets, revealing how the model aligns with dataset-specific biases in different contexts.
+A.4), more examples and analysis on the generated images (Sec.
+A.6), and additional debiasing results on less biased settings for Waterbirds plus an alternative version of BAR having known $\rho$ and bias-annotations.
 4.1 Datasets
@@
 A.1.
+Figure 3: Examples of synthetic bias-aligned images across multiple datasets.
+Each grid shows synthetic images for specific classes across biased datasets, revealing how the model aligns with dataset-specific biases in different contexts.
 4.2 Results
@@
 Supplementary Material (Sec.
-A.7) shows a histogram of the detected keywords, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced.
+A.8) shows a histogram of the detected keywords, in which it can be noted that together with a high frequency of the object class (bird), similar significant bins of the bias attributes (land, water) are also evidenced.
 4.2.2 Classification accuracy
-Table 1 reports the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered datasets.
+Tables 1(a), 1(b), 2(a), 2(b), and 3 report the average and standard deviation of our employed metrics, obtained over three independent runs for all the considered benchmark datasets.
+We dedicate a separate table for each dataset, as not all methods share the same benchmarks.
 We report Worst Group Accuracy (WGA) for Waterbirds, as in Sagawa*2020Distributionally ; Lee_Park_Kim_Lee_Choi_Choo_2023 .
@@
 It is worth noticing that, for BAR and BFFHQ, we report benchmark results related to works that employ the same versions as ours.
-Method Unsup Waterbirds BAR BFFHQ ImageNet-A
-WGA Avg. Avg. Confl. Avg.
-[…]
+Table 1: Results of DDB Recipes I and II on Waterbirds (left) and BAR (right). Ticks () mark bias-unsupervised methods. Best unsupervised results are in bold; second-best are underlined.
+(a) Debiasing results on Waterbirds
+Method Unsup Waterbirds
+WGA
+ERM $62.60\text{\scriptsize{$\pm$ 0.30}}$
 LISA yao2022improving $89.20$
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-[…]
-Park et al. park2024enhancing $71.68$
-[…]
-CDvG+LfF jung2023fighting $84.80$ $62.20\text{\scriptsize{$\pm$0.45}}$ $34.60$
-[…]
-[…]
-[…]
-[…]
-[…]
-Table 1: Results of DDB Recipes I and II on benchmark datasets. Ticks () mark bias-unsupervised methods. Best unsupervised results are in bold; second-best are underlined.
-Across all the considered benchmark datasets, both DDB recipes show state-of-the-art results (see Table 1).
+G-DRO Sagawa*2020Distributionally $91.40\text{\scriptsize{$\pm$ 1.10}}$
+George sohoni2020no $76.20\text{\scriptsize{$\pm$ 2.00}}$
+JTT pmlr-v139-liu21f_JTT $83.80\text{\scriptsize{$\pm$ 1.20}}$
+CNC pmlr-v162-zhang22z $88.50\text{\scriptsize{$\pm$ 0.30}}$
+B2T+G-DRO kim2024discovering $90.70\text{\scriptsize{$\pm$ 0.30}}$
+LfF nam2020learning $78.00$
+CDvG+LfF jung2023fighting $84.80$
+MoDAD pastore2024lookingmodeldebiasinglens $89.43\text{\scriptsize{$\pm$ 1.69}}$
+DDB-II (ours) $\bf 91.56\text{\scriptsize{$\pm$ 0.15}}$
+DDB-I (ours) $90.81\text{\scriptsize{$\pm$ 0.68}}$
+DDB-I (w/ err. set) $90.34\text{\scriptsize{$\pm$ 0.41}}$
+(b) Debiasing results on BAR
+Method Unsup BAR
+Avg.
+ERM $51.85\text{\scriptsize{$\pm$ 5.92}}$
+BiaSwap kim2021biaswap $52.40\text{\scriptsize{$\pm$ -}}$
+JTT pmlr-v139-liu21f_JTT $68.53\text{\scriptsize{$\pm$ 3.29}}$
+LfF nam2020learning $62.98\text{\scriptsize{$\pm$ 2.76}}$
+LWBC NEURIPS2022_75004615_LWBC $62.03\text{\scriptsize{$\pm$ 2.76}}$
+DebiAN li2022discover $69.88\text{\scriptsize{$\pm$ 2.92}}$
+MoDAD pastore2024lookingmodeldebiasinglens $69.83\text{\scriptsize{$\pm$ 0.72}}$
+DDB-II (ours) $\bf 72.81\text{\scriptsize{$\pm$ 1.02}}$
+DDB-I (ours) $70.40\text{\scriptsize{$\pm$ 1.41}}$
+DDB-I (w/ err. set) $70.59\text{\scriptsize{$\pm$ 0.19}}$
+Across all the considered benchmark datasets, both DDB recipes show state-of-the-art results.
 Notably, in Waterbirds, DDB’s WGA is higher than all the other state-of-the-art unsupervised methods’ accuracies, as well as of that of the supervised G-DRO Sagawa*2020Distributionally (91.56% vs.
@@
 DDB-I is providing slightly worse performance than DDB-II, but still better than the state of the art.
+Table 2: Results of DDB Recipes I and II on BFFHQ (left) and ImageNet-9/A (right) from kim2021biaswap .
+Ticks () mark bias-unsupervised methods.
+Best unsupervised results are in bold; second-best are underlined.
+(a) Debiasing results on BFFHQ
+Method Unsup BFFHQ
+Avg. Confl.
+ERM $60.13\text{\scriptsize{$\pm$ 0.46}}$
+BiaSwap kim2021biaswap $58.87\text{\scriptsize{$\pm$ -}}$
+JTT pmlr-v139-liu21f_JTT $62.20\text{\scriptsize{$\pm$ 1.34}}$
+LfF nam2020learning $62.97\text{\scriptsize{$\pm$ 3.22}}$
+ETF-Debias wang2024navigate $73.60\text{\scriptsize{$\pm$ 1.22}}$
+Park et al. park2024enhancing $71.68$
+CDvG+LfF jung2023fighting $62.20\text{\scriptsize{$\pm$0.45}}$
+DebiAN li2022discover $62.80\text{\scriptsize{$\pm$ 0.60}}$
+MoDAD pastore2024lookingmodeldebiasinglens $68.33\text{\scriptsize{$\pm$ 2.89}}$
+DDB-II (ours) $\bf 83.15\text{\scriptsize{$\pm$ 1.76}}$ $70.93\text{\scriptsize{$\pm$ 0.14}}$
+DDB-I (ours) $81.27\text{\scriptsize{$\pm$ 0.88}}$ $\bf 74.67\text{\scriptsize{$\pm$ 2.37}}$
+DDB-I (w/ err. set) $82.44\text{\scriptsize{$\pm$ 0.64}}$ $71.40\text{\scriptsize{$\pm$ 0.92}}$
+(b) Debiasing results on ImageNet-9/A
+Method Unsup ImageNet-A
+Avg.
+ERM $30.30$
+LWBC NEURIPS2022_75004615_LWBC $35.97\text{\scriptsize{$\pm$0.49}}$
+CDvG+LfF jung2023fighting $34.60$
+DDB-II (ours) $37.53\text{\scriptsize{$\pm$0.82}}$
+DDB-I (ours) $\bf 39.80\text{\scriptsize{$\pm$0.50}}$
+DDB-I (w/ err. set) $38.12\text{\scriptsize{$\pm$0.96}}$
 In BAR, DDB-II is reaching 72.81% average accuracy, which is 2.93% better than the second best (DebiAN li2022discover ), in terms of conflicting accuracy.
 In BFFHQ, our DDB-I reaches an accuracy of 74.67%, with an average improvement of 1% when compared to the state-of-the-art ETF-Debias wang2024navigate (73.60%).
+Table 3: Comparison of deep learning methods on the UrbanCars biased dataset.
+ID.ACC is in-distribution accuracy (the higher, the better); other three metrics show the accuracy gap between each minority group and ID.ACC (the lower, the better).
+GroupDRO and LLE are separated as they rely on bias information.
 Method ID.ACC $\uparrow$ BG-Gap $\downarrow$ CoObj-Gap $\downarrow$ BG-CoObj-Gap $\downarrow$
@@
 DebiAN li2022discover $\underline{98.00}$ $14.90$ $10.50$ $69.00$
-[…]
-[…]
-Table 2: Comparison of deep learning methods on the UrbanCars biased dataset.
-ID.ACC is in-distribution accuracy (the higher, the better); other three metrics show the accuracy gap between each minority group and ID.ACC (the lower, the better).
+DDB-I (ours) $86.39\text{\scriptsize{$\pm$ 0.74}}$ $\bf 1.85\text{\scriptsize{$\pm$ 3.21}}$ $\bf 0.52\text{\scriptsize{$\pm$ 1.38}}$ $\bf 0.12\text{\scriptsize{$\pm$ 1.56}}$
+DDB-II (ours) $\bf 98.56\text{\scriptsize{$\pm$ 0.92}}$ $2.30\text{\scriptsize{$\pm$ 0.60}}$ $11.10\text{\scriptsize{$\pm$ 1.20}}$ $46.70\text{\scriptsize{$\pm$ 2.42}}$
 For UrbanCars (see Table 3), presenting multiple biases, we adopted the same metrics of li2023whac (see Suppl.
@@
 WGA variations along different $w$ values equal or larger than 1 are limited, with the best performance obtained for $w=1$.
+Table 4: Ablations on: (left) synthetic image count for training the BA on Waterbirds; (right) Impact of guidance strength $w$ on CDPM sampling, evaluated on Waterbirds.
+For both experiments, we employ Recipe I.
 # Synth. images for BA training
@@
 Guidance strength
-$w$ 0 1 2 3 5
-WGA 87.85 90.81 89.25 89.56 88.32
-Table 3: Ablations on: (left) synthetic image count for training the BA on Waterbirds; (right) Impact of guidance strength $w$ on CDPM sampling, evaluated on Waterbirds, employing Recipe I.
+$w$ 0 $\mathbf{1}$ 22 33 55
+WGA $87.85$ $\mathbf{90.81}$ $89.25$ $89.56$ $88.32$
 Group extraction via BA plain error sets.
@@
 3.3.1, we describe how we filter the BA error set through a simple heuristic based on per-sample loss distribution.
-In Table 1, the last row (DDB-I w/ plain err.
+In Tables 1(a), 1(b), 2(a), and 2(b), the last row (DDB-I w/ plain err.
 set) reports our obtained results when considering as bias-conflicting all the samples misclassified by the Bias Amplifier.
 Our results show a trade-off between the two alternatives.
-The entire set of misclassifications are always enough to reach high bias mitigation performance with less critical settings (e.g., like in Waterbirds, $\rho=0.950$), gaining an advantage from such a coarser selection.
+The entire set of misclassifications is always enough to reach high bias mitigation performance with less critical settings (e.g., like in Waterbirds, $\rho=0.950$), gaining an advantage from such a coarser selection.
 On the other hand, the importance of a more precise selection is highlighted by the superior performance of the filtered alternative in the more challenging scenarios, like in BFFHQ ($\rho=0.995$).
@@
 This result shows that our Bias Amplifier can still provide beneficial signals to the target model in standard scenarios, potentially favoring the learning of the most complex samples.
+Effectiveness of Synthetic Data for BA Training.
+A core claim of our work is that training the Bias Amplifier on synthetic data prevents the memorization of real bias-conflicting samples.
+To validate this, we compare our standard BA (trained on synthetic data) with a BA trained on the original real training data within our two recipes.
+Using the real dataset significantly degrades performance.
+For Recipe I, the BA trained on real data for 50 epochs memorized all conflicting samples, leading to empty pseudo-label groups.
+Training for just one epoch (as in JTT pmlr-v139-liu21f_JTT ) yielded a WGA of only 79.43%.
+For Recipe II, using a frozen BA trained on real data resulted in a WGA of 78.45%, far below our proposed method and in line with the original LfF performance.
+These results empirically confirm that using synthetic data is crucial for creating an effective BA.
 5 Conclusions
@@
 Synthetic images sampled from the inferred biased distribution are used to train a Bias Amplifier, employed to provide a supervisory signal for training a debiased model.
-The usage of synthetic bias-aligned data prevents the detrimental effects typically impacting auxiliary models’ training in unsupervised debiasing schemes, such as bias-conflicting sample overfitting and interference, which lead to suboptimal generalization performance of the debiased target model.
-DDB results effective and versatile, acting as a plug-in for unsupervised debiasing methods.
-In this work, we incorporate DDB’s Bias Amplifier in a two-step and an end-to-end debiasing recipes.
-leading to accuracies outperforming state-of-the-art works by a margin on popular biased benchmark datasets (considering single and multiple biases), including recent works relying on vision-language models kim2024discovering ; ciranni2024say .
+The usage of synthetic bias-aligned data is empirically shown to prevent the detrimental effects typically impacting auxiliary models’ training in unsupervised debiasing schemes, such as bias-conflicting sample overfitting and interference, which lead to suboptimal generalization performance of the debiased target model.
+Our experiments confirm that a bias amplifier trained on real data fails at this task, while our approach creates a robust and effective supervisory signal.
+DDB is versatile, acting as a plug-in for unsupervised debiasing methods, since its Bias Amplifier can be incorporated into several debiasing approaches.
+In this work, we incorporate DDB’s Bias Amplifier in a two-step and an end-to-end debiasing recipes, outperforming state-of-the-art works on popular biased benchmark datasets (considering single and multiple biases), including recent works relying on vision-language models kim2024discovering ; ciranni2024say .
 Finally, while effectively mitigating bias dependency for biased datasets, DDB does not degrade performance when used with unbiased data, making it suitable to be adopted as an effective general-purpose standard training procedure.
 Limitations.
-DDB’s main limitations rely on the well-known high computational complexity of diffusion models.
-As a trade-off between computational efficiency and generation quality, we reduce the training sets resolution to $64\times 64$ when training the CDPM.
-Still, our CDPM requires $\sim 14$ hours for training on the largest considered dataset (i.e., BFFHQ, having $19,200$ images) on an NVIDIA A30 with $24$ GB of VRAM.
-Another limitation inherited from diffusion models is that the generation quality may decrease if the training dataset is too small, impacting DDB applicability for very small-scale datasets.
+DDB’s main limitations stem from the high computational cost of training diffusion models.
+As a trade-off between efficiency and generation quality, we reduce the training image resolution to $64\times 64$; nonetheless, training our CDPM still requires $\sim$14 hours on an NVIDIA A30 GPU with 24 GB of VRAM for the largest considered dataset (BFFHQ, 19,200 images).
+Our ablations show, however, that this cost can be substantially reduced by training for fewer iterations or at lower resolutions (e.g., 3.2 hours for Waterbirds at $32\times 32$) with only a minor impact on final performance.
+Importantly, this computational cost should be weighed against the significant, often impractical, burden of manually annotating a bias-aware validation set, which many alternative methods require [34, 43].
+Another limitation inherited from diffusion models is that generation quality may decrease if the training dataset is too small, potentially limiting DDB’s applicability in very small-scale settings.
+Finally, while we have benchmarked DDB on a wide range of standard datasets, the bias mitigation community currently lacks million-scale datasets with appropriate bias annotations for large-scale evaluation, which remains an important direction for future research.
+Acknowledgements.
+We acknowledge the financial support from PNRR MUR Project PE0000013 "Future Artificial Intelligence Research (FAIR)", funded by the European Union – NextGenerationEU, CUP J33C24000430007, CUP C63C22000770006.
 Appendix A Technical Appendices and Supplementary Material
@@
 It has 4,795 training images, 1,199 images for validation, and 5,794 for testing.
-[…]
+  • BFFHQ kim2021biaswap builds upon FFHQ DBLP:journals/pami/KarrasLA21 by introducing demographic biases for face recognition, while BAR nam2020learning is crafted to challenge action recognition models that associate specific actions with particular stereotypical contexts.
+For these two last datasets, we specifically refer to their original versions introduced in nam2020learning and kim2021biaswap , respectively: BFFHQ has a total of 19,200 training images where the semantic classes  $y\in\{\text{young, old}\}$, the bias attribute  $b\in\{\text{female, male}\}$, and the bias ratio  $\rho=0.995$.
+Then, it provides 1,000 validation images (all bias-aligned) and 1,000 testing images with uniformly distributed labels and bias attributes.
   • We utilize the original BAR version as introduced in nam2020learning .
@@
 Bias amplifier data requirements.
+Table 5: Ablation on the number of training images for the Bias Amplifier, in terms of WGA on Waterbirds, when using DDB’s Recipe II.
 # Synth. images for BA training
@@
 WGA $84.42$ $90.03$ $90.18$ $90.78$ $\bf 91.56$
-Table 4: Ablation on the number of training images for the Bias Amplifier, in terms of WGA on Waterbirds, when using DDB’s Recipe II.
 This ablation study measures how the number of training images for the Bias Amplifier impacts DDB’s Recipe II.
@@
 CFG strength and generation bias.
+Table 6: Impact of guidance strength $w$ on the final WGA on Waterbirds, when using DDB’s Recipe II.
 Guidance Strength
-$w$ 00 $\bf 1$ $2$ $3$ $5$
+$w$ 0 $\bf 1$ 22 33 55
 WGA $87.07$ $\bf 91.56$ $88.01$ $89.87$ $88.30$
-Table 5: Impact of guidance strength $w$ on the final WGA on Waterbirds, when using DDB’s Recipe II.
 Here, we evaluate DDB’s Recipe II dependency on the guidance strength parameter.
@@
 Also Recipe II demonstrates to be applicable in unbiased scenarios, as it measures a final test accuracy of $83.87\%$ on CIFAR10, which is slightly higher than the ERM baseline ($83.26\%$).
+DDB with less biased datasets.
+We evaluated DDB’s robustness on datasets with weaker spurious correlations by resampling Waterbirds to create versions with $\rho\in\{0.90,0.80,0.70\}$.
+Table 7 shows that while the ERM baseline improves as the bias weakens, DDB-I consistently provides a significant performance boost across all bias levels, demonstrating its effectiveness even when the bias is less pronounced.
+Table 7: WGA (%) on Waterbirds with varying bias correlation ($\rho$).
+Method $\rho=0.95$ $\rho=0.90$ $\rho=0.80$ $\rho=0.70$
+ERM 62.60 63.40 64.12 68.84
+DDB-I 90.81 86.91 90.19 86.14
+Performance on an alternative version of BAR with bias annotations.
+We tested out method on an alternative version of the BAR dataset, specifically the one introduced in Lee_Park_Kim_Lee_Choi_Choo_2023 and having $\rho=0.95$.
+Table 8 shows how even for this other version of BAR, our method is effective in mitigating bias dependency.
+When applying our Recipe-I we obtain SOTA performance, confirming the good results already obtained for the benchmark dataset in the main paper.
+Table 8: Results on the annotated BAR dataset ($\rho=0.95$).
+Method Average Accuracy
+ERM 82.40%
+ETF-Debias [46] 83.66%
+DisEnt+BE [28] 84.96%
+DDB-I (Ours) 85.36% ± 0.71
+Impact of Generated Image Quality.
+We studied the trade-off between the visual quality of generated images, computational cost, and debiasing performance on BFFHQ.
+We trained the CDPM for fewer iterations, resulting in images with higher (worse) Fréchet Inception Distance (FID) scores.
+As shown in Table 9, performance degrades gracefully as image quality decreases.
+Even with a CDPM trained for only 170k iterations (FID of 30.80), the final conflicting accuracy remains competitive with the state-of-the-art and takes only $\sim$9.5 hours to train.
+This supports our claim that photorealistic quality is secondary to capturing the essential bias cues; even images with artifacts contain the necessary signals for our BA to function effectively.
+Table 9: Impact of CDPM training iterations on FID and final conflicting accuracy (%) on BFFHQ with Recipe I.
+# CDPM Iters FID (lower is better) Conflicting Acc.
+200k (original) 22.60 74.67 ± 2.37
+170k 30.80 73.70 ± 0.71
+120k 36.39 71.20 ± 1.13
+70k 41.82 67.40 ± 2.55
 A.6 Additional Biased Generations
@@
 Figure 8: Synthetic bias-aligned image generations for each class of the BFFHQ dataset. Each grid displays 100 synthetic images per class, highlighting model bias alignment.
-A.7 Bias Validation via Captions in Synthetic Images
+A.7 Quantitative Analysis of Bias Amplification
+To empirically verify our hypothesis that a CDPM amplifies the bias present in a training dataset, we conducted a quantitative analysis on Waterbirds.
+We created training sets with varying degrees of spurious correlation ($\rho\in\{0.95,0.90,0.80,0.70\}$) and trained a separate CDPM on each.
+We then generated 1,000 synthetic images per class from each trained model.
+These synthetic datasets were manually annotated to estimate the resulting bias correlation ($\rho_{synth}$).
+As shown in Table 10, the CDPM consistently generated a dataset with a higher bias ratio than the one it was trained on.
+For example, when trained on data with $\rho=0.95$, the generated samples exhibited a correlation of $\rho_{synth}=0.99$.
+This empirically demonstrates the bias-amplifying property of diffusion models.
+This behavior can be attributed to the models learning the density of the training data and, during sampling, having a higher probability of generating samples from the high-density regions [23], which correspond to the majority bias-aligned groups in each class.
+This amplification is precisely what allows our BA to learn a robust representation of the bias without being exposed to the rare, real bias-conflicting samples.
+Table 10: Bias amplification in synthetic data generated by a CDPM trained on Waterbirds with varying training bias ratios ($\rho_{train}$).
+$\rho_{train}$ 0.950 0.900 0.800 0.700
+$\rho_{synth}$ 0.990 0.961 0.912 0.827
+A.8 Bias Validation via Captions in Synthetic Images
 As an ulterior confirmation of bias in CDPM-generated images, besides two independent annotators, we implement an unsupervised analysis pipeline using the pre-trained BLIP-2 FlanT5XL model DBLP:conf/icml/0008LSH23 for zero-shot image captioning without any task-specific prompting, for avoiding any guidance in the descriptions of the synthetic data.